### v2

# LIFBench: Evaluating the Instruction Following Performance and Stability of Large Language Models in Long-Context Scenarios

Abstract
As Large Language Models (LLMs) evolve in natural language processing (NLP), their ability to stably follow instructions in long-context inputs has become critical for real-world applications. However, existing benchmarks seldom focus on instruction-following in long-context scenarios or stability on different inputs. To bridge this gap, we introduce LIFBench, a scalable dataset designed to evaluate LLMs’ instruction-following capabilities and stability across long contexts. LIFBench comprises three long-context scenarios and eleven diverse tasks, featuring 2,766 instructions generated through an automated expansion method across three dimensions: length, expression, and variables. For evaluation, we propose LIFEval, a rubric-based assessment method that enables precise, automated scoring of complex LLM responses without reliance on LLM-assisted assessments or human judgment. This method allows for a comprehensive analysis of model performance and stability from multiple perspectives. We conduct detailed experiments on 20 prominent LLMs across six length intervals. Our work contributes LIFBench and LIFEval as robust tools for assessing LLM performance in complex and long-context settings, offering valuable insights to guide future advancements in LLM development. [^1] Data and code are available at https://github.com/SheldonWu0327/LIF-Bench-2024
LIFBench: Evaluating the Instruction Following Performance and Stability of Large Language Models in Long-Context Scenarios
Xiaodong Wu  Minhao Wang  Yichen Liu  Xiaoming Shi  He Yan §  Xiangju Lu §  Junmin Zhu §  Wei Zhang thanks: Corresponding author. East China Normal University  §iQIYI Inc {51255901079,51275901104,51275901148}@stu.ecnu.edu.cn  xmshi@ir.hit.edu.cn §{yanhe, luxiangju, zhujunmin}@qiyi.com  *zhangwei.thu2011@gmail.com
1 Introduction
  Benchmark Long. Inst. Stab. Unlim.
ZeroSCROLLS (2023)
BAMBOO (2024)
Longbench (2023)
$\infty$ Bench (2024c)
RULER (2024)
[…]
FollowBench (2023)
InfoBench (2024b)
CELLO (2024)
LIFBench (Ours)
Table 1: A comparison of our LIFBench with some relevant datasets. We summarize their focus, including Long-context scenarios, Instruction-following, and model Stability. ’Unlim.’ denotes whether the data length can be Unlimited.
Figure 1: The framework of LIFBench, where the task Single-ID in the List scenario is used as an example. Bold denotes the scenario description $D$; normal denotes the context $X$; italics denotes instruction $I$, with red indicate the instruction variables $var$, and the remaining black parts correspond to the instruction template $tpl$.
As Large Language Models (LLMs) continue to make significant strides across practical applications Achiam et al. (2023); Chowdhery et al. (2023); Brown (2020), their performance in natural language processing (NLP) tasks has reached unprecedented levels. These tasks span text generation Que et al. (2024); Tan et al. (2024); Zhang et al. (2024b), complex reasoning Parmar et al. (2024), and problem-solving Lu et al. (2024); Li et al. (2024b). Despite these achievements, significant challenges remain. On the one hand, LLMs often struggle to accurately and consistently follow human instructions, such as restating input content precisely or adhering stably to specific formatting constraints He et al. (2024). On the other hand, studies show that as input length increases, the LLMs’ performance in tasks such as reasoning Levy et al. (2024), retrieval Li et al. (2024a), and general NLP Bai et al. (2023) deteriorates. These challenges pose substantial barriers to their effectiveness in real-world applications.
Numerous evaluation benchmarks, as summarized in Table 1, have been proposed to guide the development of LLMs. However, they each exhibit notable limitations when it comes to evaluating instruction-following capabilities and stability in long-context scenarios. Some benchmarks focus either on long-context scenarios Bai et al. (2023); Zhang et al. (2024c) or complex instruction-following abilities Zhou et al. (2023a); Jiang et al. (2023). However, none of these works evaluate instruction-following abilities in long-context scenarios, and their reliance on fixed data lengths fails to accommodate the state-of-the-art LLMs’ ever-expanding context length. Other efforts Li et al. (2024a); Hsieh et al. (2024) attempt to extend evaluation to longer contexts by constructing synthetic datasets, but their tasks are hard to provide a comprehensive and in-depth assessment of LLMs’ instruction-following abilities. In terms of evaluation outcomes, most existing benchmarks focus exclusively on task completion performance, often neglecting stability—a critical factor in ensuring reliable real-world performance.
To address these limitations, we introduce the Long-context Instruction Following Benchmark (LIFBench), a scalable benchmark for evaluating LLMs’ instruction-following capability and stability in long-context scenarios. The framework of our benchmark is shown in Figure 1. Considering real-world scenarios, we construct three long-context scenarios based on the granularity of information to be processed. On this basis, eleven delicate tasks are designed, which can illustrate various dimensions of instruction-following capabilities. We manually craft templates for all tasks, and introduce an automated instruction expansion method from three dimensions (length, expression, and variables), enabling LIFBench to expand significantly in both the quantity and length of instructions. As an example, we construct a dataset of 2,766 instructions spanning six length intervals, reaching up to 128k tokens.
For evaluation, traditional metrics for downstream tasks are often unsuitable for complex instruction-following scenarios Honovich et al. (2023). Moreover, while many studies rely on GPT-4 for automated and open-ended assessment, these approaches encounter limitations due to notable gaps between GPT-4 and human experts Qin et al. (2024b); Jiang et al. (2023), as well as potential bias problems Wang et al. (2023). To address these challenges, we propose LIFEval, a systematic and efficient evaluation method for complex LLM responses, without relying on LLMs or human evaluators. Specifically, by designing task-specific scoring rubrics, we decompose evaluations into fine-grained and quantifiable scoring points, each of which can be assessed automatically. In addition, through the score-capability map and a novel metric—IFS, LIFEval provides insights into models’ fundamental capabilities and stability from various perspectives.
Overall, our contributions are as follows:
  • We introduce LIFBench, a benchmark designed to evaluate instruction-following capabilities in long-context scenarios, containing 11 tasks across three scenarios.
  • We develop methods for dataset expansion across three perspectives, enabling high scalability in both the quantity and length of instructions.
  • We propose LIFEval, an automatic evaluation method for accurately and comprehensively assessing the quality and stability of LLMs’ complex responses.
  • We conduct extensive experiments across six length intervals, which evaluate and analyze the instruction-following capabilities and stability of 20 well-known LLMs, encompassing both open-source and closed-source models.
2 Related Work
Several studies focus on LLMs’ performance in long contexts. These benchmarks collect data from traditional NLP tasks (e.g., summarization and question answering (QA)) to form comprehensive datasets containing long data Shaham et al. (2022, 2023); Dong et al. (2024); Li et al. (2024c). Additionally, due to the excellent performance of LLMs on open-ended tasks, some benchmarks have also designed synthetic tasks to better probe the models’ ability in math, reasoning, and logic Bai et al. (2023); Zhang et al. (2024c); Kwan et al. (2023); Wang et al. (2024); Hsieh et al. (2024); Li et al. (2024a). For evaluation, some works adopt regular metrics (e.g., Acc, ROUGE, and BLEU), which can be obtained through automated calculations. However, some open-ended tasks cannot be effectively evaluated using these metrics, hence powerful LLMs, such as GPT-4, are used as alternative evaluators. For example, the studies An et al. (2023); Li et al. (2023) feed the model predictions and ground-truth answers to the GPT-4 evaluator, which is tasked with conducting a direct scoring or comparison to evaluate baselines’ performance on partial tasks. In contrast to LIFBench, these benchmarks only assess problem-solving capabilities in long-context scenarios, overlooking the challenges of instruction following that arise in real-world applications.
Besides, some other efforts Jiang et al. (2023); Jing et al. (2023); Zhou et al. (2023a); Qin et al. (2024a) focus on diverse constraint types in real-world tasks, aiming to assess LLMs’ abilities in understanding and meeting these constraints. Given the complexities of evaluating instruction-following abilities, several studies introduce meaningful innovations in their evaluation methodologies to better tackle this multifaceted challenge. The studies Cook et al. (2024); Wen et al. (2024); Qin et al. (2024b); Zhang et al. (2024a) decompose instructions into checklists composed of YES/NO questions or PASS/FAIL criteria, which answers should meet. CELLO He et al. (2024) defines four criteria that should be assessed as they can encompass common errors made by LLMs, and develops automated evaluation metrics to reflect models’ ability in complex instruction-following scenarios. However, these studies fail to address model performance in long-context instruction-following scenarios and largely overlook stability in the instruction-following process.
  Scenario Task ID #Samples
List Single-ID LSI 180
Multi-ID LMI 150
Offset-ID LOI 396
Offset-Element LOE 432
Blur-ID LBI 396
Blur-Element LBE 432
MultiDoc Batch-label MB 150
Find-dup-doc MF 150
OneDoc Repeat OR 150
QA OQ 180
Extract OE 150
Table 2: Statistics of LIFBench.
3 LIFBench
3.1 Problem Definition
As shown in Figure 1, we model the instruction-following task in long-context scenarios as follows: Given a prompt consisting of a scenario description ($D$), context ($X$), and instruction ($I$), the model is expected to output an answer ($A$). This process can be represented as:
$(D,X_{len},I_{tpl,var})\to A\,.$ (1)
In this setup, the scenario description $D$ provides task background at the beginning of the prompt, and all tasks within a scenario share the same $D$. The context $X$, as the main body of the prompt, provides essential information and varies by scenario. For example, in the List scenario, $X$ is a long list, while in the OneDoc scenario, $X$ represents a lengthy processed document. The parameter $len$ represents the number of tokens in $X$. The instruction $I$ is placed at the end of the prompt, which consists of two components: (1) the instruction template ($tpl$), which outlines the task requirements, and (2) the instruction variable ($var$), representing variable part within the template. Generally, in LIFBench, $D$ and $I$ tend to be short, while $X$ is a long text with thousands of tokens.
3.2 Dataset Construction
To accurately simulate real-world application scenarios for LLMs in long-text processing, we construct 3 scenarios and 11 corresponding tasks (see Table 2), enabling a comprehensive and precise evaluation of instruction-following capabilities.
List   The List scenario is designed to assess the ability of LLMs to handle structured batches of short-item information, as seen in applications like retrieval, recommendation, and structured data processing. The context $X$ is a lengthy ordered list, and LLMs must identify elements that meet the specified criteria provided in the instructions.
The scenario includes six tasks: Single-ID and Multi-ID focus on retrieving specific elements from an ordered list using provided IDs. Building on these basic tasks, the Offset-ID, Offset-element, Blur-ID, and Blur-element tasks introduce spatial constraints, adding complexity to the retrieval process. "Offset" tasks require precise index-based retrieval to test fine-grained spatial awareness, while "Blur" tasks involve broader spatial ranges, allowing more flexibility. "ID" and "Element" refer to the type of reference in the retrieval process, representing either the position number in the ordered list or the list element itself.
MultiDoc   The MultiDoc scenario tests LLMs on tasks involving coarser-grained information, such as multi-document summarization, clustering, and retrieval-augmented generation (RAG). Unlike the List scenario, the MultiDoc tasks require models to compare multiple documents, identify differences, and execute batch operations, demanding holistic attention to the input rather than focusing on a few key elements.
To construct context $X$, we curate and process datasets from diverse domains, combining them into multi-document collections with high diversity. There are two tasks in this scenario. Find-dup-doc requires LLMs to identify and output multiple duplicate documents in a specified format. In batch-label, models label each document based on specified attributes.
OneDoc   The OneDoc scenario simulates tasks involving processing of a single and very long document, typically occurring in use cases like information extraction and question answering. In this scenario, LLMs are required to identify and process the most relevant information from a vast amount of text, testing their ability to resist noise and focus on task-specific instructions.
To create this scenario, we construct an extra-long document by concatenating entries from the Paul Graham Essays dataset [^2] https://huggingface.co/datasets/sgoel9/paul_graham_essays, and tagging some sentences as key information. Then three tasks are designed: The Repeat task requires models to repeat a specified amount of key information. The Extract task instructs models to extract specific types of key information. The QA task presents models with a sentence and asks it to identify whether the sentence is labeled as key information.
We manually write scenario descriptions and instruction templates for all tasks, with detailed examples provided in Appendix H. To further ensure the tasks and scenarios are challenging and discriminative, special efforts are made during data collection, as elaborated in Appendix A.
3.3 Data Extension
In this section, we expand manual templates in three dimensions (length, expression, and variable) to form a sizeable test dataset.
Length   A number of works Bai et al. (2023); Ni et al. (2024); Li et al. (2024c); Levy et al. (2024) have found that the length of input text has an impact on the ability of the LLMs. In LIFBench, different lengths of prompts allow us to explore the impact of context length on the instruction-following capabilities of LLMs, making it essential to introduce variations in prompt length.
In all three scenarios, we adjust the length of the prompt by controlling the context token count $l$. Specifically, we modify the number of elements in the List or the number of documents in MultiDoc and OneDoc to achieve the desired length. We have pre-constructed ample corpus for each scenario, supporting expansions up to 2M tokens in one prompt.
Expression   In real-world contexts, due to personality and individual differences, individuals often provide significantly varied descriptions of the same subject, which undoubtedly introduces challenges to the stability of large models in following instructions. To assess LLM robustness in this regard, we diversify instruction templates to create multiple expressions style with differing wording and syntax.
Our approach follows a four-step process, namely "Rewriting-Encoding-Clustering-Sampling (RECS)". First, we use GPT-4 Achiam et al. (2023) and Claude Anthropic (2024) to generate 40+ rewrites of each original instruction template. Next, the rewritten templates are encoded into vector representations for clustering, with the number of clusters set to the target number of rewritten instructions for each task. For instance, if a task needs five rewritten instructions, we will create five clusters. Finally, we sample from each cluster, selecting the usable template closest to the cluster center as the final diversified expression. More implementation details and evidence of effectiveness can be found in Appendix B.
Variable   Some placeholders are preset in the instruction templates to indicate variable parts of instructions, i.e., instruction variables. These variables encompass elements such as query keys (for retrieval tasks), categorization criteria (for classification tasks), and format requirements. By analyzing LLM performance across varying instruction variables, we can evaluate the models’ understanding and consistency in executing instructions. Ideally, models with strong instruction-following abilities execute instructions stably across variable conditions, while weaker models may exhibit inconsistency. Each instruction variable has been manually assigned by specific sampling spaces and rules, enabling the generation of a comprehensive instruction set. For specific examples, please refer to Appendix C.
Table 2 details the task counts for each scenario. Compared to the other two scenarios, tasks in the List scenario are simpler, so they carry less weight (see Section 4.1) but are more numerous.
Figure 2: Rubric score distribution across different scenarios and capabilities. We carefully designed the rubric to ensure scores are as evenly distributed as possible across different scenarios and capabilities.
4 LIFEval
In this section, we introduce LIFEval, an automatic method that provides accurate assessments and detailed insights into LLMs’ long-context instruction-following capabilities and stability.
4.1 Automated Rubric-based Scoring
To evaluate the output quality of LLMs on LIFBench tasks, we introduce Automated Rubric-based Scoring (ARS), an accurate and efficient programmatic evaluation method. As shown in Figure 1 and Figure 2, we manually craft scoring rubrics $\mathcal{R}$ for each task. Each rubric consists of several scoring points $s$ and is assigned a weight $\tilde{s}$ according to its importance and complexity. In other words, a larger $\tilde{s}$ means a more complex task, requiring more evaluative consideration. All of these points can be assessed automatically through a program. The scoring process on task $t$ can be shown as follows:
$\begin{split}ARS_{t}(\mathcal{A}_{t})=\frac{1}{\tilde{R_{t}}}\sum_{s\in\mathcal{R}_{t}}f_{s}(\mathcal{A}_{t}),\tilde{R}_{t}=\sum_{s\in\mathcal{R}_{t}}\tilde{s}\,.\end{split}$ (2)
In this equation, $\mathcal{A}_{t}$ represents the model’s responses on task $t$, and $\tilde{R}_{t}$ is the sum of weights across all scoring points. The function $f_{s}(\cdot)\in[0,\tilde{s}]$ represents the average score of all outputs for scoring point $s$, as determined by the automated evaluation program. Naturally, defining $\mathcal{T}=\{t_{i}\}_{i=1}^{N_{t}}$ as the set of all tasks, the overall test result for any model on LIFBench is the weighted average of scores across $\mathcal{T}$:
$ARS_{overall}=\frac{\sum_{t\in\mathcal{T}}\tilde{R}_{t}\cdot ARS_{t}}{\sum_{t\in\mathcal{T}}\tilde{R}_{t}}\,.$ (3)
  OneDoc List MultiDoc
Model OR OQ OE LSI LMI LOI LOE LBI LBE MB MF Overall
API Models
[…]
GPT-4 0.707 0.820 0.736 0.893 0.735 0.704 0.781 0.750 0.832 0.600 0.777 0.738
Models Larger Than 20B Parameters
[…]
Llama-3.1-70B-Inst. 0.730 0.860 0.711 0.805 0.798 0.651 0.798 0.693 0.788 0.657 0.531 0.694
Qwen2.5-32B-Inst.$\dagger$ 0.702 0.792 0.850 0.662 0.612 0.542 0.517 0.763 0.761 0.608 0.476 0.650
C4AI-cmd-r-08-2024 (32B)$\dagger$ 0.529 0.838 0.535 0.692 0.619 0.494 0.530 0.667 0.615 0.729 0.660 0.626
C4AI-cmd-r-v01 (35B)$\dagger$ 0.495 0.818 0.420 0.641 0.579 0.489 0.506 0.694 0.643 0.721 0.646 0.595
[…]
Llama-3.1-70B 0.337 0.273 0.118 0.668 0.394 0.525 0.581 0.695 0.681 0.708 0.308 0.422
[…]
Models With 7-20B Parameters
[…]
InternLM2.5-7b-chat-1m 0.446 0.828 0.378 0.609 0.438 0.543 0.631 0.713 0.764 0.619 0.428 0.523
Qwen2.5-7B-Inst.$\dagger$ 0.507 0.812 0.447 0.626 0.568 0.436 0.531 0.684 0.701 0.445 0.436 0.519
Llama-3.1-8B-Inst. 0.537 0.681 0.413 0.705 0.535 0.397 0.537 0.668 0.637 0.522 0.308 0.491
GLM-4-9b-chat-1m 0.484 0.813 0.267 0.705 0.534 0.371 0.514 0.648 0.667 0.688 0.300 0.490
[…]
LWM-Text-Chat-1M (7B) 0.413 0.730 0.075 0.633 0.291 0.309 0.605 0.590 0.590 0.128 0.520 0.381
Llama-3.1-8B 0.347 0.287 0.040 0.600 0.207 0.422 0.554 0.625 0.622 0.471 0.455 0.375
[…]
LWM-Text-1M (7B) 0.307 0.252 0.049 0.164 0.136 0.110 0.306 0.420 0.452 0.112 0.220 0.204
Table 3: The ARS scores of models on different tasks. The abbreviations of the tasks can be found in Table 2. The overall score is calculated by Eq. 3. The best performing score is highlighted in bold and second-best is underlined. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
4.2 Score-Capability Mapping
To further offer insights into the model’s strengths and weaknesses across various dimensions of instruction following, we introduce Score-Capability Mapping, which maps the scoring point $s$ to six fundamental capabilities. With reference to previous studies He et al. (2024); Zhou et al. (2023b), the six capabilities are defined as follows:
Ori   (Original Content): Abilities to reproduce the original input accurately.
Num   (Numerical Ability): Abilities in handling numerical data, such as recognition, counting, and basic arithmetic.
Spat   (Spatial Awareness): Abilities in understanding spatial relationships and sequences.
Fmt   (Format): Abilities in modifying and structuring content according to format rules.
Logic   (Logic Execution): Abilities to follow logical conditions and decision branches.
Recog   (Recognition Ability): Abilities to differentiate and focus on key elements of the input.
Building on the ARS score, we further compute the Instruction Following Performance (IFP) for each capability. The IFP for a specific capability $c$ is defined as:
$IFP_{c}=\frac{\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{R}_{t}}\mathbb{I}(s,c)\cdot f_{s}(\mathcal{A})}{\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{R}_{t}}\mathbb{I}(s,c)\cdot\tilde{s}}\,$ (4)
where $\mathbb{I}(s,c)$ is the indicator function that identifies whether a scoring point $s$ is related to the capability $c$. We manually crafte this mapping so that when a scoring point $s$ is relevant to the capability $c$, the indicator $\mathbb{I}(s,c)$ equals 1, and 0 otherwise.
4.3 Instruction Following Stability
To measure whether the model can consistently follow instructions, we introduce Instruction Following Stability (IFS). Specifically, we define the observation perspective of stability, $p$, which refers to a specific feature of the model input. In LIFBench, there are three perspectives: prompt length, expression (i.e., the template of instruction $I$), and instruction variables. Based on the selected perspective, we partition model’s responses $\mathcal{A}_{t}$ into $N_{p}$ groups, denoted as $\mathcal{A}_{t}^{p}=\{\mathcal{A}_{t}^{j}\}_{j=0}^{N_{p}}$. For example, a set of responses $\mathcal{A}_{t}$ may originate from inputs spanning five different length intervals (such as 4k, 8k, etc.). In this case, we group the responses into five distinct groups based on the input lengths. Subsequently, ARS scores for each group of answers will be computed, resulting in a set of performance values $\mathcal{Y}_{t}^{p}$, which can be expressed as follows:
$\begin{split}\mathcal{Y}_{t}^{p}=\{ARS_{t}(\mathcal{A}_{t}^{j})\}_{j=0}^{N_{p}}\,.\end{split}$ (5)
Finally, the Instruction Following Stability (IFS) on task $t$ is calculated as the variance (denoted by $\sigma(\cdot)$) of the performance values $Y$ divided by their mean, formally expressed as:
$\begin{split}IFS_{t}^{p}=\frac{\sigma(\mathcal{Y}_{t}^{p})}{\bar{\mathcal{Y}}_{t}^{p}}\in[0,+\infty)\end{split}$ (6)
A lower IFS indicates greater stability in instruction-following under perspective $p$, whereas a higher IFS signifies reduced stability.
5 Experiments
5.1 Experiment Setup
We evaluate 20 popular LLMs with long-context capabilities, including models from GPT Achiam et al. (2023), Llama Dubey et al. (2024), Qwen Yang et al. (2024), C4AI Cohere For AI (2024), LWM Liu et al. (2024a), InternLM Cai et al. (2024), and GLM GLM et al. (2024) series, all claiming to support context lengths exceeding 128k tokens. Notably, the Qwen2.5-Inst. model extends its 32k context length to 128k with YaRN Peng et al. (2023). GPT-4 and GPT-4o are accessed via its official API, while open-source models are deployed using vLLM Kwon et al. (2023).
Experiments are conducted across six context length intervals from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation. Token counts are calculated using GPT-4’s tokenizer [^3] https://github.com/openai/tiktoken, and truncation is applied to adjust context $X$ for models unable to process the longest contexts. Further details about the experimental setup can be found in the Appendix F.
  IFS Overall
Model Expression Variable Length IFS(AVG) ARS
API Models
[…]
GPT-4 0.066 (2) 0.101 (8) 0.155 (3) 0.107 (3) 0.738 (2)
Models Larger Than 20B Parameters
[…]
C4AI-cmd-r-v01$\dagger$ 0.135 (9) 0.082 (5) 0.143 (2) 0.120 (4) 0.596 (7)
Qwen2.5-32B-Inst.$\dagger$ 0.103 (5) 0.087 (6) 0.182 (6) 0.124 (5) 0.651 (5)
Llama-3.1-70B-Inst. 0.101 (4) 0.076 (2) 0.263 (13) 0.147 (8) 0.694 (4)
C4AI-cmd-r-08-2024$\dagger$ 0.114 (7) 0.077 (4) 0.238 (11) 0.143 (7) 0.626 (6)
[…]
Llama-3.1-70B 0.225 (18) 0.165 (19) 0.380 (17) 0.257 (17) 0.433 (14)
Models With 7-20B Parameters
[…]
Qwen2.5-7B-Inst.$\dagger$ 0.142 (10) 0.112 (9) 0.250 (12) 0.168 (13) 0.520 (11)
Qwen2.5-14B-Inst.$\dagger$ 0.133 (8) 0.124 (15) 0.206 (9) 0.154 (9) 0.548 (9)
Llama-3.1-8B-Inst. 0.154 (13) 0.120 (12) 0.199 (8) 0.158 (12) 0.491 (12)
GLM-4-9b-chat-1m 0.151 (12) 0.138 (18) 0.175 (5) 0.155 (10) 0.490 (13)
Llama-3.1-8B 0.215 (17) 0.119 (11) 0.319 (16) 0.217 (15) 0.392 (17)
[…]
LWM-Text-Chat-1M 0.204 (16) 0.130 (17) 0.282 (15) 0.206 (14) 0.411 (15)
LWM-Text-1M 0.237 (19) 0.179 (20) 0.280 (14) 0.232 (16) 0.205 (20)
[…]
Table 4: Instruction Following Stability from different perspectives. The integers in parentheses represent rankings, where smaller numbers indicate better performance. The overall order in each partition depends on the average IFS rankings from three perspectives. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
5.2 Results on LIFBench
Task-categorized Performance   As shown in Table 3, two closed-source models achieve the highest scores, with the top score reaching only 0.758. This highlights substantial room for improvement in instruction-following capabilities. Generally, larger parameter sizes are associated with better performance. However, models with supervised instruction tuning often outperform their base counterparts, even when smaller in size. For instance, InternLM2.5-7b-chat-1m scores 0.101 higher than Llama-3.1-70B, demonstrating that fine-tuning on instruction or conversational data significantly enhances performance. While closed-source models dominate most tasks, open-source models excel in only a few (e.g., Blur-ID and Batch-label). Overall, open-source small and medium models still lag behind closed-source large models.
Capability-categorized Performance   As shown in Figure 3, closed-source models lead in all dimensions. Within each model series, the solid lines consistently enclose their corresponding dashed lines, highlighting the significant impact of instruction fine-tuning on overall performance. Notably, the Format dimension exhibits the most compact clustering, suggesting that formatting requirements are well-covered across training corpora for most models. In contrast, the Recog dimension shows marked differences, likely reflecting varying levels of effort in incorporating numerical cognition data during training.
Stability   Table 4 presents the IFS scores and rankings from three perspectives, revealing discrepancies between model stability and task completion ability. For instance, while Qwen2.5-72B-Inst. underwent truncation, potentially affecting its stability in "Length", it still outperformed GPT-4, which had a higher ARS score. Furthermore, models exhibited distinct strengths and weaknesses across perspectives: GPT-4o showed less stability in "Expression", while GPT-4 struggled with instruction variables where Llama-3.1-70B-Inst. excelled. Instruction fine-tuning generally improved stability , but larger parameter size did not guarantee better performance, as Qwen2.5-72B-Inst. surpassed all closed-source models in "Expression".
Figure 3: The Instruction Following Performance in six capabilities. Lines of the same color in the chart represent models from the same series. Dashed lines represent base models, while solid lines represent their officially fine-tuned variants.
Figure 4: Overall ARS score (a) and instruction following stability (b) under different input length.
Figure 5: Instruction following performance in six core capabilities under different input length.
5.3 Effects of longer inputs
Overall Performance and Stability   As shown in Figure 4(a), the performance of most models declines significantly as input length increases, particularly beyond 16k or 32k tokens. However, the rates of decline vary across models. For instance, GPT-4o maintains relatively high scores in long-context scenarios, whereas Llama-3-70B-Inst. experiences a sharp drop, indicating its limited ability to handle extended inputs.
The negative impact of input length is also evident in stability metrics. As illustrated in Figure 4(b), most models demonstrate poorer stability in long-context scenarios, with the lowest stability observed at the longest input lengths. Interestingly, some models, such as C4AI-cmd-r-08-2024 and GLM-4-9b-chat-1m, exhibit the least stability at mid-range input lengths (e.g., 64k tokens), diverging from their overall performance trends. Additionally, the sensitivity to input length also varies across different perspectives. For instance, models like C4AI-cmd-r-08-2024 and Qwen2.5-7B-Inst. show more significant upward trend in "Expression" compared to "Variable". This highlights potential areas for improvement in enhancing instruction-following stability under long input contexts.
Performance in six core capabilities   Figure 5 demonstrates the declining trends of six core capabilities across varying model sizes as input length increases. Notably, "Format" performance remains relatively stable across all input lengths in most models, suggesting that tasks related to formatting are less sensitive to longer contexts. Conversely, "Recog" experiences the steepest decline, highlighting the challenges models face in maintaining recognition ability as input length grows.
Model size also plays a crucial role in long-context instruction-following performance. Larger models generally are better at capabilities like "Ori," "Num," and "Spat," indicating that handling original content and spatial reasoning demands more model parameters. In contrast, format-related abilities are effectively managed even by smaller models, suggesting they are less dependent on model size.
6 Conclusion
In this work, we systematically evaluate LLMs’ instruction-following capabilities and stability in long-context scenarios. We develop LIFBench, a benchmark that encompasses three long-context scenarios and a diverse set of 11 tasks, complemented by a method for instruction expansion across three distinct perspectives. For evaluation, we introduce LIFEval, an automated rubric-based scoring method, enabling fast and accurate assessments of model task performance and stability. Based on the benchmark and scoring method, we conduct extensive experiments on 20 prominent LLMs, revealing significant room for improvement in instruction-following capabilities and stability, especially in long-context scenarios.
Limitation
Our study has several limitations. Firstly, due to constraints in programmatic validation, our task scenarios lack comprehensive support for semantic constraints, necessitating future improvements for better validation. Secondly, the inference process for very long inputs requires significant computational resources and time, limiting the scale of our dataset and potentially affecting the reproducibility of our work. As a result, our benchmark example comprises fewer than 3,000 samples, tested across only three perspectives. In fact, many unexplored aspects of stability remain, such as LLMs’ consistency in handling input formatting and context domain shifts. Additionally, larger datasets can enable more statistically significant conclusions. We encourage the community to use our proposed protocol to expand the evaluation set and conduct more extensive analyzes.
Finally, while LIFEval enables efficient and automated evaluations, the reliability of its results depends heavily on the design of the scoring rubric and the implementation of the evaluation programs. These components require significant time and effort to prepare before conducting evaluations. We leave the automation or reduction of these manual efforts for future work.
Appendix A Details in Data Collection
Figure 6: IFS Gap Across Three Perspectives with RECS vs. Random Sampling. Positive values indicate an IFS increase with RECS compared to Random Sampling, while negative values indicate a decrease. "MEAN" on the x-axis represents the average IFS gap.
A.1 List
To ensure the quality of tasks, the context $X$ for the List scenario is constructed as a ordered list, consisting of two types of elements: randomly generated UUIDs and natural language instruction texts. Liu et al. (2024b) used UUIDs to build key-value pairs and explained how large models utilize input context. Inspired by this work, we generated a set of unique 128-bit UUIDs as the first part of the list. However, real-world scenarios often involve some level of semantic noise, and transformer-based language models may exhibit varying sensitivities to different linguistic features in their input O’Connor and Andreas (2021). Therefore, to enhance complexity and realism, we selected a subset of instruction texts from the Alpaca-52k dataset Taori et al. (2023), which serve as the second type of elements in the list. We found that when instruction texts are mixed into the retrieval list, the model’s attention tends to be drawn to the embedded instructions, leading to a prioritization of following them rather than focusing on the originally assigned task. As a result, we chose the "Instructions" part of the Alpaca-52k dataset as list elements. In addition, to ensure the appropriateness of text length, all selected instructions were limited to 5~40 tokens.
A.2 MultiDoc
To construct contexts $X$ in this scenario, we selected documents from four datasets: GovReport Huang et al. (2021), XSum Narayan et al. (2018), QMSum Zhong et al. (2021), and the Paul Graham Essays [^4] https://huggingface.co/datasets/sgoel9/paul_graham_essays. GovReport is a long document summary dataset consisting of reports written by government research agencies; XSum and QMSum are respectively a news summary dataset and a metting summary dataset , both of which cover a wide range of domains; and the Paul Graham Essays dataset collects articles written by Paul Graham in a variety of fields. These selective datasets span multiple domains and various forms of text, offering a high degree of diversity.
For each dataset, we extract the main text (excluding summaries) to construct multi-document task contexts $X$, limiting each text to 300–500 tokens through filtering and truncation. As shown in Table LABEL:tab:allprompts, each document contains six fields: "text", "id", "iD2", "title", "date", and "source". The "text" field holds the processed content, while the other five are attributes designed for the tasks, partially sourced from the dataset and partially constructed manually. Each document has unique "id" and "iD2" fields, though some may lack "title" or "source". To reduce LLMs’ reliance on parameter knowledge, we randomly annotate the "source" field, breaking its correlation with the "text". This ensures the model cannot infer the source using pre-trained knowledge, creating additional challenges for instruction adherence. Lastly, we introduced duplicates by reusing some "text" fields in different documents, maintaining a duplication rate of 25%.
A.3 OneDoc
To create the context $X$ for OneDoc, we synthesized an extra-long document by concatenating entries from the Paul Graham Essays dataset, following the approach in Greg Kamradt (2023). Additionally, some sentences are randomly tagged as key information, with each tag specifying a type (e.g., Topic, Evidence, Concession) and a unique identifier to aid LLMs in identifying and categorizing critical content.
Appendix B Effectiveness of Expression Extension
Implementation details   To implement RECS for expression extension, we chose the BGE-M3 text embedding model Chen et al. (2024) for encoding and applied K-means clustering Macqueen (1967). Additionally, since LLMs can introduce inaccuracies in rewrites due to misinterpretations, we manually filtered out unsuitable rewrites during the sampling phase, resulting in 5~10 instruction templates for each task.
Prompt for Rewriting in RECS Please rewrite the given prompt according to the following requirements: 1. The rewritten prompt must retain the same meaning as the original, without altering the intent. 2. Try your best to use different vocabulary or sentence structures. 3. Ensure that the rewritten prompt is clear and accurate, avoiding any expressions that could lead to ambiguity. 4. Please keep the placeholders in the prompt (i.e., “{}” and the contents therein) exactly as they are during rewriting. 5. Please keep the example in the prompt, but you can make some small changes while keeping the original meaning. 6. Output the result in Json List format, without anything else. 7. Please generate 20 different rewrites at once. prompt: {Prompt to be rewritten}
Figure 7: Prompt template for rewriting task prompt.
Validity Experiment   To validate the effectiveness of the RECS method proposed in Section 3.3, we compared it with a random sampling approach. Specifically, after the Rewriting phase, we constructed a dataset by randomly sampling (rather than clustering) the same number of instructions and calculated the IFS values for sixteen models on this dataset.
As shown in 6, compared to random sampling, RECS led to higher expression IFS scores in 75% of the models, outperforming the IFS improvement rates observed for the Variable and Length perspectives, at 37.5% and 62.5% respectively. Additionally, when comparing the mean IFS values across all models, we observed that RECS improved expression IFS while slightly lowering variable IFS and length IFS. This suggests that the RECS method make it more challenging for models to maintain stability when handling varied expressions, which shows the effectiveness in expression extension as well.
Appendix C Sampling Space for Instruction Variables
As shown in Figure 8, we reserved identical placeholders for each instruction template and set the sampling space accordingly. Depending on the task requirements, the types of instruction variables include numerical values, lists, phrases, long sentences, and format indicators, etc.
List: Single-ID Instruction templates: 1. Retrieve the entry at position {pos} in the list. Display it immediately. 2. What is the {pos} element in the list? Output it directly. 3. … Sampling Space
pos: [1st, 3rd, 8th, …]
Instruction after sampling
1. Retrieve the entry at position 8th in the list. Display it immediately.
2. …
OneDoc: QA Instruction templates:
1. Is the phrase "{evidence}" the KEY SENTENCE? Reply with "{options[0]}" for yes and "{options[1]}" for no.
2. Is "{evidence}" the KEY SENTENCE of the text? Indicate "{options[0]}" for yes, "{options[1]}" for no. 3. … Sampling Space
evidence: [a sentence sampled from the input document.]
options: [(True, False), (Yes, No), (apple, banana), …] Instruction after sampling
1.Is "With the result that …" the KEY SENTENCE of the text? Indicate "True" for yes, "False" for no.
2. …
Figure 8: Examples of the sampling space. Red indicates the instruction variables $var$, and {} indicates a placeholder in the instruction.
Appendix D Rubric Design and Program Evaluation
Rubric Design   As shown in Table 5 and Figure 2, we manually designed a scoring rubric for LIFBench, ensuring a balanced evaluation across different scenarios and capabilities. The rubric assigns greater weight $\tilde{s}$ to score point $s$ that is more challenging. For instance, in assessing format correctness, compared to the QA (OQ) task, the Batch-Label (MB) task are assigned higher scoring weights due to more complex formatting requirements.
Program Evaluation   As described in Section 4.1, we design an automated evaluation program based on the scoring rubric to evaluate the quality of LLM responses. Inspired by Zhou et al. (2023a), during task design, we aim to ensure that the correct answers could be captured by the program. However, our tasks incorporate more complex formatting and logical constraints, making it challenging to directly obtain accurate evaluation results through simple automated methods. To address this, we iteratively decomposed and refined each scoring point $s$ into sub-evaluation criteria that could be directly assessed by the program, enabling a more detailed and discriminative evaluation.
Our program evaluation adheres to two principles: (1) correct answers should achieve full scores, and (2) partially correct answers should be distinguishable in terms of scores. In general, we first extract structured information from the outputs based on the format constraints of each task, and then perform evaluations on other dimensions using this information. However, LLM responses are not always well-structured and may contain minor errors that hinder the extraction of structured information. To mitigate this, we carefully designed task-specific validation programs that employ techniques such as regex matching and substring retrieval to maximize the extraction of valid part from LLM responses. Finally, to ensure evaluation quality, we manually provided reference answers for all examples in LIFBench, verifying that they can achieve full scores in our automated evaluation process.
Appendix E Inspiration for IFS
IFS is derived from the coefficient of variation (CV) Brown (1998), a statistical measure widely used in fields such as risk assessment and quality control Shechtman (2013). CV is calculated as the ratio of the standard deviation to the mean, providing a normalized measure of variability.
Using standard deviation alone to measure instruction-following stability can lead to misleading conclusions. For instance, if a model’s performance across all intervals is zero, its standard deviation would also be zero, suggesting perfect stability despite the model’s complete inability to perform. IFS avoids this issue by normalizing variability relative to the mean, ensuring that stability is assessed in a scale-independent and interpretable manner.
The application of CV to instruction-following tasks aligns well with its traditional use in evaluating consistency and reliability across varying scenarios. By capturing fluctuations in model performance relative to its average capability, IFS offers a robust and fair metric for comparing the stability of different LLMs.
Appendix F Experiments Setup
Baselines   For closed-source models, we specifically use gpt-4o-2024-08-06 and gpt-4-0125-preview to represent the performance of GPT-4o and GPT-4, respectively. For open-source models, we test the performance of both base models and fine-tuned models. Models with the suffix "-Inst." or "-chat" indicate that they have been fine-tuned on instruction or dialogue data.
Inference   During the inference process, we complete the deployment of all open-source models with the vLLM Kwon et al. (2023). The temperature is set to 0 to ensure deterministic outputs. For the SFT versions of the models, we used the official chat template. For the base versions, we add a suffix "Output: " to prompt the models to generate answers according to the instructions. In order to avoid unnecessary time consumption resulting from endless repetitions in the output, we set maximum generation lengths for different tasks: List scenario tasks that output a single element were limited to 100 tokens, tasks in the MultiDoc scenario are set to 4096 tokens, and the rest of the tasks are limited to 512 tokens.
Due to variations in encoding efficiency across models’ tokenizer, some models can not accommodate the longest inputs. To address this, we apply right truncation to the context $X$, ensuring the completeness of scenario description $D$ and instruction $I$.
  Tasks Score Point $s$ Weight $\tilde{s}$ Related Capability
Format correctness. 1 Fmt
Answer is from the input list. 2 Ori
Answer correctness. 1 Recog
[…]
Format correctness. 2 Fmt
Order correctness. 2 Spat
Number of output elements is correct. 3 Num
Answers correctness. 3 Ori
[…]
Format correctness. 1 Fmt
Answer is from the input list. 1 Ori
The output element conform to the position constraint. 3 Spat
[…]
Answer correctness (conforms to label logic). 3 Logit
Output labels are from the candidate set. 3 Ori
The number of output labels matches the number of input documents. 3 Num, Recog
Format correctness. 5 Fmt
[…]
Answer correctness. 4 Logit, Recog
Find the correct number of duplicate documents. 5 Num, Logit
Format correctness. 5 Fmt
Document properties in the output are in the input 6 Ori
[…]
Answer correctness. 3 Logit
Output sentences are from the input document. 2 Ori
Format correctness. 3 Fmt
Output sentences are the key sentence. 2 Recog
The number of output key sentences is correct. 4 Num
[…]
Answer correctness. 3 Logit
Format correctness. 2 Fmt
[…]
Output sentences are from the input document. 2 Ori
Format correctness. 4 Fmt
Output sentences are target key sentences. 4 Recog
The output order matches the ids sort. 4 Spat
[…]
Table 5: The scoring rubric and Score-Capability Map in LIFEval.
Appendix G Full Results on LIFBench
G.1 ARS Score
  OneDoc List MultiDoc
Model Length OR OQ OE LSI LMI LOI LOE LBI LBE MB MF Overall
4k 0.770 0.831 0.781 0.942 0.938 0.835 0.874 0.755 0.866 0.735 0.627 0.776
8k 0.859 0.863 0.893 0.942 0.946 0.707 0.856 0.754 0.830 0.782 0.637 0.807
16k 0.830 0.863 0.812 0.925 0.885 0.778 0.849 0.794 0.856 0.749 0.710 0.801
32k 0.799 0.901 0.831 0.863 0.871 0.745 0.808 0.767 0.871 0.778 0.617 0.779
64k 0.758 0.935 0.829 0.829 0.755 0.773 0.826 0.728 0.753 0.746 0.468 0.721
GPT-4o 128k 0.766 0.895 0.857 0.783 0.619 0.601 0.726 0.699 0.775 0.526 0.468 0.666
4k 0.939 0.832 0.907 0.875 0.923 0.789 0.819 0.790 0.896 0.702 0.879 0.859
8k 0.953 0.865 0.926 0.954 0.871 0.795 0.811 0.790 0.869 0.715 0.836 0.854
16k 0.777 0.793 0.802 0.954 0.801 0.769 0.824 0.731 0.885 0.744 0.843 0.804
32k 0.527 0.787 0.750 0.975 0.684 0.715 0.807 0.746 0.815 0.552 0.756 0.700
64k 0.526 0.819 0.537 0.833 0.584 0.648 0.761 0.687 0.808 0.681 0.663 0.647
GPT-4 128k 0.524 0.825 0.494 0.767 0.547 0.510 0.665 0.757 0.719 0.207 0.685 0.561
4k 0.823 0.773 0.907 0.954 0.853 0.767 0.813 0.820 0.867 0.698 0.876 0.833
8k 0.840 0.781 0.892 0.942 0.782 0.774 0.731 0.858 0.846 0.697 0.783 0.805
16k 0.722 0.727 0.822 0.900 0.630 0.703 0.682 0.837 0.781 0.704 0.738 0.741
32k 0.768 0.731 0.803 0.842 0.615 0.714 0.741 0.876 0.841 0.684 0.557 0.710
64k 0.814 0.874 0.685 0.850 0.589 0.583 0.594 0.755 0.812 0.713 0.332 0.645
Qwen2.5-72B-Inst.$\dagger$ 128k 0.583 0.759 0.789 0.717 0.574 0.535 0.524 0.765 0.693 0.157 0.220 0.502
4k 0.952 0.838 0.941 0.983 0.898 0.803 0.863 0.842 0.831 0.555 0.709 0.814
8k 0.884 0.864 0.943 0.979 0.948 0.809 0.919 0.800 0.884 0.765 0.702 0.842
16k 0.903 0.806 0.894 0.892 0.884 0.786 0.858 0.802 0.841 0.778 0.617 0.804
32k 0.560 0.843 0.829 0.875 0.897 0.680 0.809 0.726 0.813 0.630 0.568 0.707
64k 0.609 0.928 0.624 0.742 0.761 0.525 0.770 0.715 0.831 0.772 0.381 0.645
Llama-3.1-70B-Inst. 128k 0.475 0.882 0.036 0.358 0.402 0.302 0.571 0.270 0.531 0.444 0.209 0.353
4k 0.920 0.767 0.897 0.736 0.706 0.626 0.587 0.789 0.826 0.840 0.800 0.808
8k 0.728 0.855 0.919 0.692 0.719 0.560 0.548 0.840 0.796 0.826 0.633 0.749
16k 0.670 0.754 0.907 0.721 0.574 0.565 0.547 0.789 0.756 0.711 0.529 0.678
32k 0.736 0.815 0.858 0.683 0.596 0.572 0.602 0.748 0.719 0.146 0.402 0.578
64k 0.598 0.809 0.829 0.646 0.540 0.549 0.475 0.731 0.780 0.742 0.255 0.597
Qwen2.5-32B-Inst.$\dagger$ 128k 0.557 0.750 0.687 0.494 0.540 0.380 0.341 0.682 0.689 0.386 0.238 0.489
4k 0.620 0.849 0.819 0.783 0.767 0.738 0.718 0.750 0.809 0.773 0.936 0.791
8k 0.689 0.824 0.682 0.742 0.758 0.569 0.659 0.744 0.838 0.827 0.824 0.755
16k 0.609 0.869 0.573 0.858 0.588 0.420 0.531 0.646 0.484 0.786 0.752 0.664
32k 0.499 0.895 0.546 0.763 0.570 0.471 0.564 0.648 0.531 0.803 0.592 0.616
64k 0.391 0.769 0.360 0.608 0.477 0.401 0.400 0.602 0.500 0.809 0.454 0.512
C4AI-cmd-r-08-2024 128k 0.364 0.822 0.228 0.400 0.556 0.366 0.308 0.612 0.530 0.373 0.402 0.416
4k 0.664 0.798 0.559 0.746 0.650 0.536 0.505 0.692 0.604 0.799 0.916 0.715
8k 0.570 0.863 0.450 0.767 0.668 0.520 0.502 0.685 0.566 0.765 0.794 0.659
16k 0.496 0.797 0.386 0.567 0.573 0.528 0.470 0.637 0.598 0.711 0.636 0.577
32k 0.461 0.842 0.448 0.588 0.555 0.442 0.558 0.690 0.676 0.832 0.593 0.598
64k 0.360 0.780 0.356 0.663 0.512 0.495 0.574 0.731 0.780 0.763 0.467 0.541
C4AI-cmd-r-v01 128k 0.417 0.829 0.321 0.517 0.515 0.414 0.429 0.729 0.633 0.458 0.471 0.482
4k 0.576 0.738 0.674 0.709 0.734 0.560 0.604 0.697 0.779 0.815 0.834 0.722
8k 0.505 0.753 0.598 0.775 0.611 0.516 0.503 0.671 0.726 0.765 0.691 0.647
16k 0.360 0.705 0.500 0.592 0.641 0.529 0.504 0.681 0.648 0.760 0.500 0.563
32k 0.294 0.647 0.484 0.425 0.504 0.426 0.504 0.663 0.728 0.839 0.501 0.538
64k 0.321 0.596 0.073 0.508 0.462 0.362 0.523 0.597 0.598 0.891 0.291 0.434
[…]
4k 0.889 0.187 0.095 0.767 0.635 0.686 0.631 0.778 0.745 0.431 0.295 0.494
8k 0.091 0.264 0.047 0.633 0.423 0.530 0.553 0.606 0.694 0.915 0.367 0.414
16k 0.441 0.290 0.179 0.783 0.446 0.547 0.509 0.721 0.644 0.892 0.221 0.461
3k2 0.117 0.300 0.330 0.713 0.361 0.598 0.686 0.715 0.728 0.899 0.308 0.458
64k 0.218 0.300 0.021 0.742 0.282 0.603 0.713 0.716 0.729 0.852 0.179 0.390
Llama-3.1-70B 128k 0.264 0.300 0.036 0.370 0.219 0.190 0.394 0.633 0.547 0.261 0.479 0.311
4k 0.515 0.809 0.553 0.586 0.554 0.487 0.570 0.586 0.579 0.792 0.508 0.588
8k 0.582 0.813 0.483 0.561 0.585 0.444 0.530 0.687 0.651 0.757 0.567 0.602
16k 0.338 0.817 0.429 0.055 0.306 0.187 0.368 0.254 0.221 0.859 0.307 0.413
32k 0.296 0.503 0.144 0.033 0.100 0.005 0.000 0.052 0.000 0.926 0.397 0.313
64k 0.128 0.133 0.036 0.297 0.358 0.223 0.250 0.255 0.196 0.907 0.310 0.311
[…]
4k 0.843 0.761 0.731 0.697 0.740 0.480 0.365 0.787 0.736 0.796 0.534 0.695
8k 0.630 0.780 0.751 0.583 0.791 0.454 0.446 0.798 0.715 0.687 0.497 0.649
16k 0.601 0.748 0.680 0.440 0.566 0.547 0.372 0.667 0.506 0.739 0.367 0.569
32k 0.501 0.812 0.609 0.603 0.481 0.483 0.377 0.702 0.475 0.438 0.285 0.485
64k 0.518 0.766 0.568 0.484 0.519 0.472 0.435 0.670 0.524 0.617 0.211 0.491
Qwen2.5-14B-Inst.$\dagger$ 128k 0.463 0.740 0.484 0.341 0.510 0.308 0.318 0.577 0.468 0.269 0.203 0.394
4k 0.608 0.809 0.492 0.688 0.455 0.691 0.751 0.712 0.834 0.734 0.682 0.648
8k 0.490 0.869 0.540 0.675 0.384 0.592 0.682 0.678 0.868 0.744 0.625 0.617
16k 0.399 0.841 0.439 0.738 0.434 0.530 0.637 0.717 0.727 0.710 0.338 0.523
32k 0.386 0.847 0.329 0.638 0.449 0.594 0.704 0.683 0.803 0.767 0.334 0.518
64k 0.435 0.821 0.213 0.488 0.489 0.475 0.551 0.762 0.721 0.227 0.313 0.414
InternLM2.5-7b-chat-1m 128k 0.361 0.781 0.256 0.429 0.416 0.377 0.459 0.727 0.630 0.534 0.275 0.419
4k 0.545 0.850 0.543 0.692 0.615 0.693 0.660 0.808 0.745 0.637 0.723 0.656
8k 0.484 0.831 0.569 0.717 0.623 0.493 0.611 0.783 0.744 0.679 0.604 0.623
16k 0.516 0.815 0.621 0.717 0.556 0.454 0.589 0.672 0.618 0.573 0.506 0.578
32k 0.555 0.866 0.335 0.713 0.530 0.563 0.612 0.722 0.710 0.112 0.278 0.444
64k 0.483 0.787 0.321 0.575 0.554 0.245 0.372 0.500 0.693 0.510 0.274 0.445
Qwen2.5-7B-Inst.$\dagger$ 128k 0.462 0.724 0.292 0.346 0.531 0.166 0.340 0.619 0.698 0.156 0.230 0.366
  OneDoc List MultiDoc
Model Length OR OQ OE LSI LMI LOI LOE LBI LBE MB MF Overall
4k 0.713 0.609 0.532 0.721 0.680 0.437 0.607 0.739 0.540 0.569 0.436 0.580
8k 0.547 0.734 0.353 0.754 0.668 0.392 0.545 0.639 0.599 0.630 0.372 0.527
16k 0.624 0.679 0.517 0.783 0.536 0.467 0.613 0.731 0.731 0.638 0.292 0.548
32k 0.426 0.648 0.167 0.642 0.530 0.399 0.557 0.646 0.655 0.623 0.340 0.457
64k 0.464 0.543 0.552 0.725 0.428 0.403 0.530 0.703 0.688 0.356 0.219 0.446
Llama-3.1-8B-Inst. 128k 0.449 0.874 0.355 0.604 0.366 0.287 0.373 0.548 0.612 0.317 0.191 0.388
4k 0.578 0.837 0.388 0.738 0.692 0.425 0.633 0.702 0.749 0.743 0.441 0.589
8k 0.450 0.877 0.262 0.788 0.616 0.415 0.511 0.671 0.747 0.742 0.431 0.540
16k 0.498 0.746 0.322 0.825 0.499 0.428 0.463 0.623 0.613 0.795 0.299 0.509
32k 0.503 0.751 0.212 0.717 0.468 0.375 0.555 0.670 0.661 0.784 0.252 0.482
64k 0.423 0.830 0.166 0.688 0.484 0.314 0.467 0.646 0.604 0.713 0.196 0.437
GLM-4-9b-chat-1m 128k 0.452 0.837 0.254 0.475 0.443 0.272 0.458 0.574 0.626 0.354 0.183 0.383
4k 0.301 0.719 0.489 0.708 0.602 0.467 0.348 0.503 0.401 0.783 0.455 0.519
8k 0.260 0.673 0.472 0.548 0.527 0.360 0.459 0.516 0.197 0.730 0.383 0.462
16k 0.248 0.613 0.476 0.525 0.356 0.409 0.423 0.545 0.389 0.801 0.660 0.518
32k 0.202 0.397 0.036 0.131 0.198 0.207 0.224 0.252 0.407 0.851 0.366 0.324
64k 0.303 0.453 0.036 0.100 0.172 0.180 0.227 0.285 0.276 0.829 0.357 0.326
[…]
4k 0.373 0.644 0.036 0.654 0.331 0.478 0.631 0.722 0.615 0.252 0.519 0.403
8k 0.399 0.718 0.036 0.692 0.318 0.181 0.528 0.650 0.560 0.069 0.530 0.364
16k 0.426 0.721 0.261 0.638 0.264 0.141 0.595 0.432 0.571 0.129 0.548 0.395
32k 0.442 0.740 0.036 0.671 0.361 0.562 0.643 0.717 0.597 0.142 0.520 0.409
64k 0.405 0.790 0.044 0.629 0.231 0.292 0.656 0.591 0.639 0.087 0.502 0.366
LWM-Text-Chat-1M 128k 0.435 0.770 0.036 0.517 0.243 0.198 0.576 0.426 0.558 0.087 0.503 0.346
4k 0.473 0.251 0.014 0.733 0.361 0.681 0.747 0.773 0.722 0.788 0.635 0.521
8k 0.270 0.273 0.069 0.733 0.231 0.604 0.611 0.679 0.613 0.777 0.507 0.441
16k 0.415 0.287 0.017 0.753 0.151 0.447 0.455 0.660 0.607 0.705 0.404 0.403
32k 0.357 0.267 0.057 0.529 0.142 0.344 0.562 0.636 0.652 0.166 0.388 0.311
64k 0.276 0.340 0.056 0.513 0.114 0.413 0.532 0.588 0.600 0.188 0.381 0.298
Llama-3.1-8B 128k 0.290 0.307 0.030 0.340 0.243 0.042 0.414 0.414 0.541 0.200 0.415 0.278
4k 0.192 0.624 0.450 0.033 0.100 0.000 0.026 0.000 0.000 0.363 0.388 0.264
8k 0.214 0.665 0.353 0.084 0.103 0.022 0.103 0.020 0.073 0.633 0.233 0.274
16k 0.245 0.301 0.438 0.314 0.113 0.226 0.311 0.393 0.299 0.000 0.183 0.229
32k 0.323 0.470 0.036 0.000 0.110 0.013 0.013 0.016 0.007 0.371 0.219 0.184
64k 0.298 0.565 0.087 0.125 0.136 0.118 0.125 0.229 0.334 0.075 0.209 0.193
[…]
4k 0.334 0.293 0.036 0.259 0.126 0.233 0.311 0.565 0.428 0.151 0.230 0.230
8k 0.315 0.187 0.069 0.240 0.109 0.081 0.359 0.466 0.471 0.151 0.241 0.220
16k 0.279 0.240 0.050 0.116 0.125 0.081 0.335 0.307 0.450 0.113 0.194 0.186
32k 0.325 0.261 0.036 0.148 0.139 0.149 0.280 0.482 0.414 0.111 0.266 0.216
64k 0.294 0.267 0.059 0.178 0.113 0.118 0.304 0.512 0.519 0.128 0.190 0.208
LWM-Text-1M 128k 0.293 0.267 0.044 0.042 0.205 0.000 0.250 0.189 0.429 0.016 0.197 0.167
Table 6: ARS scores in different input lengths $l$. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
G.2 Instruction Following Performance in Six Capabilities
  Models Format Logit Num Ori Recog Spat
GPT-4o 0.712 0.582 0.842 0.836 0.756 0.806
GPT-4 0.789 0.509 0.765 0.921 0.615 0.729
Qwen2.5-72B-Inst.$\dagger$ 0.727 0.468 0.703 0.840 0.640 0.718
Llama-3.1-70B-Inst. 0.766 0.509 0.684 0.747 0.626 0.699
Qwen2.5-32B-Inst.$\dagger$ 0.760 0.400 0.671 0.752 0.492 0.647
C4AI-cmd-r-08-2024 (32B)$\dagger$ 0.721 0.480 0.678 0.833 0.453 0.508
C4AI-cmd-r-v01 (35B)$\dagger$ 0.767 0.400 0.596 0.811 0.379 0.470
[…]
Llama-3.1-70B 0.492 0.253 0.437 0.594 0.290 0.410
[…]
Qwen2.5-14B-Inst.$\dagger$ 0.659 0.400 0.612 0.591 0.410 0.523
InternLM2.5-7b-chat-1m 0.727 0.401 0.576 0.644 0.275 0.438
Qwen2.5-7B-Inst.$\dagger$ 0.707 0.359 0.612 0.595 0.317 0.420
Llama-3.1-8B-Inst. 0.465 0.349 0.539 0.627 0.393 0.553
GLM-4-9b-chat-1m 0.672 0.363 0.604 0.548 0.329 0.364
[…]
LWM-Text-Chat-1M 0.505 0.265 0.370 0.578 0.186 0.312
Llama-3.1-8B 0.586 0.328 0.331 0.521 0.088 0.303
[…]
LWM-Text-1M 0.293 0.105 0.103 0.374 0.044 0.208
Table 7: Instruction following performance in six core capabilities. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
  Models Length Format Logit Num Ori Recog Spat Models Length Format Logit Num Ori Recog Spat
4k 0.683 0.519 0.833 0.851 0.896 0.853 4k 0.786 0.653 0.827 0.999 0.900 0.889
8k 0.712 0.630 0.859 0.852 0.909 0.826 8k 0.809 0.701 0.850 0.997 0.840 0.860
16k 0.715 0.619 0.873 0.906 0.836 0.846 16k 0.771 0.545 0.851 0.983 0.792 0.815
32k 0.732 0.633 0.887 0.866 0.741 0.834 32k 0.813 0.425 0.744 0.885 0.526 0.723
64k 0.737 0.592 0.783 0.792 0.630 0.773 64k 0.839 0.373 0.666 0.896 0.377 0.564
GPT-4o 128k 0.693 0.502 0.816 0.751 0.520 0.706 GPT-4 128k 0.719 0.358 0.654 0.768 0.257 0.524
4k 0.786 0.564 0.817 0.994 0.883 0.837 4k 0.793 0.661 0.762 0.852 0.831 0.840
8k 0.798 0.545 0.761 0.953 0.806 0.818 8k 0.791 0.680 0.857 0.889 0.899 0.891
16k 0.771 0.424 0.708 0.962 0.682 0.720 16k 0.808 0.650 0.790 0.882 0.761 0.848
32k 0.741 0.437 0.724 0.892 0.584 0.734 32k 0.798 0.365 0.632 0.829 0.610 0.780
64k 0.667 0.562 0.653 0.741 0.517 0.618 64k 0.784 0.474 0.592 0.745 0.463 0.627
Qwen2.5-72B-Inst.$\dagger$ 128k 0.599 0.276 0.555 0.503 0.369 0.581 Llama-3.1-70B-Inst. 128k 0.625 0.223 0.470 0.283 0.193 0.206
4k 0.893 0.548 0.793 0.977 0.710 0.715 4k 0.844 0.659 0.868 0.959 0.707 0.727
8k 0.864 0.402 0.740 0.898 0.643 0.704 8k 0.823 0.661 0.882 0.879 0.622 0.655
16k 0.793 0.409 0.693 0.842 0.507 0.652 16k 0.729 0.481 0.666 0.959 0.522 0.476
32k 0.643 0.375 0.599 0.628 0.425 0.642 32k 0.729 0.458 0.617 0.863 0.400 0.491
64k 0.706 0.406 0.615 0.673 0.399 0.648 64k 0.631 0.361 0.533 0.775 0.291 0.379
Qwen2.5-32B-Inst.$\dagger$ 128k 0.662 0.261 0.583 0.495 0.268 0.520 C4AI-cmd-r-08-2024$\dagger$ 128k 0.573 0.260 0.503 0.562 0.175 0.321
4k 0.759 0.625 0.767 0.944 0.606 0.555 4k 0.692 0.465 0.699 0.964 0.685 0.727
8k 0.754 0.505 0.730 0.891 0.515 0.501 8k 0.629 0.322 0.509 0.928 0.643 0.652
16k 0.742 0.376 0.553 0.825 0.358 0.462 16k 0.577 0.394 0.561 0.813 0.421 0.596
32k 0.797 0.353 0.562 0.856 0.331 0.469 32k 0.613 0.323 0.485 0.847 0.318 0.533
64k 0.779 0.302 0.477 0.742 0.273 0.464 64k 0.630 0.318 0.477 0.578 0.219 0.308
[…]
4k 0.457 0.353 0.558 0.589 0.430 0.502 4k 0.498 0.477 0.679 0.771 0.550 0.616
8k 0.447 0.258 0.452 0.618 0.342 0.363 8k 0.560 0.459 0.713 0.839 0.448 0.615
16k 0.549 0.250 0.512 0.644 0.340 0.405 16k 0.457 0.440 0.491 0.508 0.305 0.385
32k 0.538 0.268 0.452 0.643 0.308 0.512 32k 0.432 0.345 0.367 0.456 0.241 0.124
64k 0.489 0.203 0.385 0.570 0.251 0.394 64k 0.419 0.243 0.448 0.476 0.190 0.138
[…]
4k 0.733 0.628 0.767 0.729 0.633 0.622 4k 0.815 0.583 0.746 0.811 0.417 0.542
8k 0.717 0.527 0.677 0.689 0.554 0.665 8k 0.813 0.550 0.710 0.753 0.373 0.520
16k 0.676 0.415 0.595 0.647 0.462 0.527 16k 0.771 0.350 0.553 0.618 0.290 0.424
32k 0.649 0.293 0.554 0.501 0.339 0.458 32k 0.734 0.355 0.554 0.648 0.264 0.439
64k 0.653 0.296 0.565 0.540 0.291 0.492 64k 0.633 0.277 0.414 0.505 0.116 0.353
Qwen2.5-14B-Inst.$\dagger$ 128k 0.527 0.240 0.513 0.404 0.180 0.377 InternLM2.5-7b-chat-1m 128k 0.594 0.289 0.480 0.531 0.190 0.352
4k 0.841 0.487 0.692 0.810 0.470 0.536 4k 0.413 0.457 0.605 0.737 0.590 0.685
8k 0.813 0.465 0.752 0.759 0.399 0.503 8k 0.460 0.465 0.628 0.623 0.503 0.545
16k 0.749 0.381 0.628 0.728 0.401 0.456 16k 0.519 0.373 0.563 0.680 0.471 0.623
32k 0.644 0.307 0.506 0.426 0.247 0.387 32k 0.476 0.332 0.549 0.686 0.274 0.423
64k 0.652 0.270 0.600 0.486 0.245 0.321 64k 0.436 0.252 0.460 0.577 0.291 0.623
Qwen2.5-7B-Inst.$\dagger$ 128k 0.545 0.242 0.496 0.361 0.141 0.317 Llama-3.1-8B-Inst. 128k 0.485 0.217 0.429 0.456 0.230 0.418
4k 0.726 0.482 0.727 0.629 0.461 0.488 4k 0.413 0.394 0.533 0.739 0.553 0.500
8k 0.716 0.458 0.682 0.579 0.423 0.374 8k 0.394 0.390 0.528 0.667 0.432 0.476
16k 0.713 0.358 0.624 0.605 0.341 0.332 16k 0.491 0.363 0.459 0.886 0.395 0.504
32k 0.672 0.323 0.591 0.584 0.286 0.350 32k 0.367 0.322 0.405 0.561 0.206 0.192
64k 0.626 0.312 0.522 0.498 0.261 0.321 64k 0.525 0.275 0.335 0.458 0.182 0.158
[…]
4k 0.640 0.313 0.370 0.534 0.107 0.334 4k 0.570 0.425 0.640 0.784 0.379 0.408
8k 0.557 0.329 0.345 0.506 0.081 0.251 8k 0.555 0.331 0.497 0.688 0.299 0.322
16k 0.570 0.365 0.351 0.548 0.087 0.361 16k 0.495 0.325 0.451 0.645 0.253 0.278
32k 0.625 0.333 0.386 0.543 0.103 0.331 32k 0.477 0.178 0.250 0.464 0.065 0.317
64k 0.573 0.318 0.258 0.517 0.083 0.293 64k 0.467 0.167 0.176 0.461 0.069 0.311
LWM-Text-Chat-1M 128k 0.554 0.311 0.275 0.480 0.069 0.245 Llama-3.1-8B 128k 0.463 0.161 0.205 0.423 0.051 0.236
4k 0.244 0.389 0.351 0.290 0.236 0.254 4k 0.317 0.145 0.138 0.424 0.045 0.233
8k 0.307 0.384 0.389 0.350 0.200 0.236 8k 0.288 0.034 0.034 0.469 0.061 0.249
16k 0.183 0.161 0.129 0.428 0.070 0.404 16k 0.280 0.099 0.095 0.325 0.042 0.176
32k 0.334 0.247 0.250 0.185 0.114 0.066 32k 0.300 0.135 0.129 0.417 0.041 0.196
64k 0.314 0.209 0.146 0.233 0.047 0.182 64k 0.310 0.102 0.077 0.362 0.034 0.238
[…]
Table 8: Instruction following performance in different input lengths $l$. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
G.3 Instruction Following Stability
  Models 4k 8k 16k 32k 64k 128k
GPT-4o 0.102 0.124 0.100 0.124 0.140 0.155
GPT-4 0.073 0.075 0.086 0.116 0.121 0.188
Qwen2.5-72B-Inst.$\dagger$ 0.100 0.103 0.092 0.094 0.120 0.136
Llama-3.1-70B-Inst. 0.127 0.124 0.118 0.154 0.150 0.261
Qwen2.5-32B-Inst.$\dagger$ 0.109 0.137 0.152 0.164 0.152 0.169
C4AI-cmd-r-08-2024$\dagger$ 0.121 0.133 0.177 0.180 0.229 0.226
C4AI-cmd-r-v01$\dagger$ 0.171 0.191 0.233 0.185 0.191 0.189
[…]
Llama-3.1-70B 0.485 0.510 0.301 0.311 0.301 0.235
[…]
Qwen2.5-14B-Inst.$\dagger$ 0.150 0.184 0.181 0.207 0.187 0.196
InternLM2.5-7b-chat-1m 0.145 0.157 0.152 0.173 0.202 0.265
Qwen2.5-7B-Inst.$\dagger$ 0.153 0.146 0.181 0.217 0.259 0.296
Llama-3.1-8B-Inst. 0.221 0.244 0.203 0.271 0.228 0.225
GLM-4-9b-chat-1m 0.165 0.163 0.195 0.214 0.254 0.229
[…]
LWM-Text-Chat-1M 0.164 0.341 0.334 0.196 0.373 0.379
Llama-3.1-8B 0.356 0.349 0.276 0.330 0.322 0.406
[…]
LWM-Text-1M 0.200 0.366 0.469 0.333 0.349 0.608
Table 9: Instruction following stability (Expression) in different input lengths $l$. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
  Models 4k 8k 16k 32k 64k 128k
GPT-4o 0.093 0.085 0.101 0.105 0.117 0.164
GPT-4 0.089 0.097 0.119 0.145 0.183 0.187
Qwen2.5-72B-Inst.$\dagger$ 0.074 0.108 0.132 0.141 0.179 0.142
Llama-3.1-70B-Inst. 0.113 0.075 0.125 0.122 0.198 0.273
Qwen2.5-32B-Inst.$\dagger$ 0.136 0.132 0.177 0.165 0.152 0.184
C4AI-cmd-r-08-2024$\dagger$ 0.140 0.156 0.148 0.156 0.205 0.147
C4AI-cmd-r-v01$\dagger$ 0.167 0.146 0.151 0.184 0.175 0.196
[…]
Llama-3.1-70B 0.405 0.500 0.272 0.183 0.172 0.220
[…]
Qwen2.5-14B-Inst.$\dagger$ 0.162 0.186 0.193 0.161 0.192 0.145
InternLM2.5-7b-chat-1m 0.191 0.155 0.181 0.217 0.336 0.240
Qwen2.5-7B-Inst.$\dagger$ 0.156 0.169 0.204 0.216 0.252 0.187
Llama-3.1-8B-Inst. 0.149 0.167 0.183 0.320 0.212 0.249
GLM-4-9b-chat-1m 0.205 0.195 0.187 0.227 0.230 0.239
[…]
LWM-Text-Chat-1M 0.158 0.159 0.256 0.195 0.234 0.206
Llama-3.1-8B 0.384 0.233 0.268 0.211 0.235 0.427
[…]
LWM-Text-1M 0.158 0.379 0.331 0.281 0.311 0.533
Table 10: Instruction following stability (Variable) in different input lengths $l$. $\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
Appendix H Examples in LIFBench
Table 11: Examples of all tasks in LIFBench. Bold denotes the scenario description $D$; normal denotes the context $X$; italics denotes instruction $I$, with red indicate the instruction variables $var$, and the remaining black parts correspond to the instruction template $tpl$.
Task Name Test Example
Scenario: List
Task: Single-ID
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction: Retrieve the entry at position 8th in the list. Display it immediately.
Scenario: List
Task: Multi-ID
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction: Identify the items at the corresponding places in list [1, 5, 7] and deliver the result in JSON list form.
Scenario: List
Task: Offset-ID
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction: Please identify the next item to 8th in the list described above.
Scenario: List
Task: Offset-Element
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction: "Considering the arrangement of the list, what is the next element after "4f78d6f47f5111ef8b42581122bf941e"?
Scenario: List
Task: Blur-ID
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction:Randomly select an item from the list above, after the 8th element.
Scenario: List
Task: Blur-Element
You’re a searcher. You need to output the corresponding list elements based on the instructions and the list below. Please follow the instructions directly without anything else.
List to be retrieved:
1. 4f63efbe7f5111ef8b42581122bf941e
2.Summarize the Rome Statute of the International Criminal Court.
153. 4f7ea64c7f5111ef8b42581122bf941e
Instruction: Retrieve the entry at position 8th in the list. Display it immediately.
Scenario: MultiDoc
Task: Find-dup-doc
You are a document manager. Here is a collection of documents. Each document includes information such as title, date, source, id, iD2 and specific article content (text). You need to read the documents and follow the instructions to give some information directly, without something else. Also note:
1. Some documents may be missing information such as title or source, which may affect the final output.
2. Some article conent (i.e. values corresponding to the text keyword) may be duplicated.
Documents:
******************** doc-1 ********************
text: Following discussions with the PSNI, … \̈newline source: meeting
iD2: bec60ab1-35bc-417d-b378-0b2e86dfcd23
id: xWTp7xEyQrqGmGj5p6CdVQ
title: 39272756
date: 2016-06-20
******************** doc-2 ********************
id: SOl-GT2FSyKFYJzBhkYzxw
text: What did the participants think about using CD’s for backup? …
source: news
date: 2001-01-15
iD2: 12a0f93d-0f57-46bc-bf28-9ec856fc974a
title: tr-sq-167
******************** doc-3 ********************
Instruction: Within the supplied documents, certain documents contain replicated content in their ’text’ field, although other fields (such as id, iD2, date, title, source) may be different. Additionally, there could be N sets of documents, each set comprising any number of replicates. Please identify these replicated documents and present iD2 in sequence. The output should have N lines, each line symbolizing a set of replicated documents. Format the output for each document set as depicted in the example. Avoid providing explanations. If a document lacks information in a specific field, use ’None’ instead.
output example:
[["iD2_1"], ["iD2_2"], ["iD2_3"]]
[["iD2_4"], ["iD2_5"]]
Scenario: MultiDoc
Task: Batch-label
You are a document manager. Here is a collection of documents. Each document includes information such as title, date, source, id, iD2 and specific article content (text). You need to read the documents and follow the instructions to give some information directly, without something else. Also note:
1. Some documents may be missing information such as title or source, which may affect the final output.
2. Some article conent (i.e. values corresponding to the text keyword) may be duplicated.
Documents:
******************** doc-1 ********************
text: Following discussions with the PSNI, … \̈newline source: meeting
iD2: bec60ab1-35bc-417d-b378-0b2e86dfcd23
id: xWTp7xEyQrqGmGj5p6CdVQ
title: 39272756
date: 2016-06-20
******************** doc-2 ********************
id: SOl-GT2FSyKFYJzBhkYzxw
text: What did the participants think about using CD’s for backup? …
source: news
date: 2001-01-15
iD2: 12a0f93d-0f57-46bc-bf28-9ec856fc974a
title: tr-sq-167
******************** doc-3 ********************
Instruction: Assign labels to documents in order using the provided list of [’11311’, ’22422’, ’33233’, ’44444’]. The labeling rules to follow are:
1. If the document contains both title and source information, mark it as "11311".
2. If the document is missing the source information but not the title, mark it as "22422".
3. If the document is missing title information but not source, mark it as "33233".
4. If the document is missing both title and source information, mark it as "44444".
The tags should be output in JSON dictionary format, for example:
{"doc1":"11311","doc2":"22422"}
Scenario: OneDoc
Task: Repeat
There are several different types of KEY SENTENCE in the input text, which are marked by special tags. These special tags a total of six kinds, respectively is <#Topic#>, <@argument@>, "<!Transition!>", "<|Summary|>", "<*Evidence*>", "<-Concession->". Different tags represent different types of key sentence. If a sentence in the text is KEY SENTENCE, we will add a special tag with the same attribute to the beginning and end of the sentence. The head tag also contains id order information in the format <type-id>. For example, the head tag with type ’#Topic#’ and id 1 is <#Topic#-1>. Also note that when the head tag and tail tag attributes are inconsistent, this means that the sentence is a fake KEY SENTENCE. Please read the input text carefully and give the answer directly according to the instruction requirements.
Input text: Remember the essays you had to write in high school? Topic sentence, introductory paragraph, supporting paragraphs, conclusion. <#Topic#-2>With the result that writing is made to seem boring and pointless.<#Topic#> Who cares about symbolism in Dickens? …
Instruction: Provide 4 KEY SENTENCE and their categories directly. Display the output on 4 individual lines with each line containing a KEY SENTENCE and its category, separated by |||||.
Output example:
[KEY_SENTENCE_1] ||||| #Topic#
[KEY_SENTENCE_2] ||||| *Evidence*
Scenario: OneDoc
Task: Extract
There are several different types of KEY SENTENCE in the input text, which are marked by special tags. These special tags a total of six kinds, respectively is <#Topic#>, <@argument@>, "<!Transition!>", "<|Summary|>", "<*Evidence*>", "<-Concession->". Different tags represent different types of key sentence. If a sentence in the text is KEY SENTENCE, we will add a special tag with the same attribute to the beginning and end of the sentence. The head tag also contains id order information in the format <type-id>. For example, the head tag with type ’#Topic#’ and id 1 is <#Topic#-1>. Also note that when the head tag and tail tag attributes are inconsistent, this means that the sentence is a fake KEY SENTENCE. Please read the input text carefully and give the answer directly according to the instruction requirements.
Input text: Remember the essays you had to write in high school? Topic sentence, introductory paragraph, supporting paragraphs, conclusion. <#Topic#-2>With the result that writing is made to seem boring and pointless.<#Topic#> Who cares about symbolism in Dickens? …
Instruction: Gather every instance of KEY SENTENCE classified as #Topic#. The output should be a Json list arranged by ids. If none are found, provide an empty array.
Output Example 1: [KEY SENTENCE1, KEY SENTENCE2, KEY SENTENCE3,…]
Output Example 2: []
Scenario: OneDoc
Task: QA
There are several different types of KEY SENTENCE in the input text, which are marked by special tags. These special tags a total of six kinds, respectively is <#Topic#>, <@argument@>, "<!Transition!>", "<|Summary|>", "<*Evidence*>", "<-Concession->". Different tags represent different types of key sentence. If a sentence in the text is KEY SENTENCE, we will add a special tag with the same attribute to the beginning and end of the sentence. The head tag also contains id order information in the format <type-id>. For example, the head tag with type ’#Topic#’ and id 1 is <#Topic#-1>. Also note that when the head tag and tail tag attributes are inconsistent, this means that the sentence is a fake KEY SENTENCE. Please read the input text carefully and give the answer directly according to the instruction requirements.
Input text: Remember the essays you had to write in high school? Topic sentence, introductory paragraph, supporting paragraphs, conclusion. <#Topic#-2>With the result that writing is made to seem boring and pointless.<#Topic#> Who cares about symbolism in Dickens? …
Instruction: Is "With the result that writing is made to seem boring and pointless." the KEY SENTENCE of the text? Indicate "True" for yes, "False" for no.

### v3

@@
 LIFBench: Evaluating the Instruction Following Performance and Stability of Large Language Models in Long-Context Scenarios
-Xiaodong Wu  Minhao Wang  Yichen Liu  Xiaoming Shi  He Yan §  Xiangju Lu §  Junmin Zhu §  Wei Zhang thanks: Corresponding author.
-East China Normal University  §iQIYI Inc {51255901079,51275901104,51275901148}@stu.ecnu.edu.cn  xmshi@ir.hit.edu.cn §{yanhe, luxiangju, zhujunmin}@qiyi.com  *zhangwei.thu2011@gmail.com
+Xiaodong Wu  Minhao Wang  Yichen Liu  Xiaoming Shi  He Yan §  Xiangju Lu §  Junmin Zhu §  Wei Zhang †‡thanks: Corresponding author.
+East China Normal University  §iQIYI Inc  Shanghai Innovation Institute {51255901079,51275901104,51275901148}@stu.ecnu.edu.cn  xmshi@ir.hit.edu.cn §{yanhe, luxiangju, zhujunmin}@qiyi.com  *zhangwei.thu2011@gmail.com
 1 Introduction
@@
 FollowBench (2023)
-InfoBench (2024b)
+InfoBench (2024)
 CELLO (2024)
@@
 (2024b), complex reasoning Parmar et al.
+(2024); Chen et al.
 (2025), and problem-solving Lu et al.
@@
 Moreover, while many studies rely on GPT-4 for automated and open-ended assessment, these approaches encounter limitations due to notable gaps between GPT-4 and human experts Qin et al.
-(2024b); Jiang et al.
+(2024); Jiang et al.
 (2023), as well as potential bias problems Wang et al.
@@
 (2024); Li et al.
-(2024c).
-Additionally, due to the excellent performance of LLMs on open-ended tasks, some benchmarks have also designed synthetic tasks to better probe the models’ ability in math, reasoning, and logic Bai et al.
+(2024c); Gavin et al.
+(2024).
+Additionally, due to the excellent performance of LLMs on open-ended tasks, some benchmarks have also designed synthetic tasks to better probe the models’ ability in math, reasoning, and logic Kwan et al.
 (2023); Zhang et al.
-(2024c); Kwan et al.
-(2023); Wang et al.
-(2024); Hsieh et al.
+(2024c); Wang et al.
 (2024); Li et al.
-(2024a).
-For evaluation, some works adopt regular metrics (e.g., Acc, ROUGE, and BLEU), which can be obtained through automated calculations.
+(2024a); Chen et al.
+(2024b).
+For evaluation, some works adopt regular metrics (e.g., Acc, ROUGE, and BLEU), which can be obtained by automated calculations.
 However, some open-ended tasks cannot be effectively evaluated using these metrics, hence powerful LLMs, such as GPT-4, are used as alternative evaluators.
@@
 (2023) feed the model predictions and ground-truth answers to the GPT-4 evaluator, which is tasked with conducting a direct scoring or comparison to evaluate baselines’ performance on partial tasks.
-In contrast to LIFBench, these benchmarks only assess problem-solving capabilities in long-context scenarios, overlooking the challenges of instruction following that arise in real-world applications.
-Besides, some other efforts Jiang et al.
-(2023); Jing et al.
-(2023); Zhou et al.
-(2023a); Qin et al.
-(2024a) focus on diverse constraint types in real-world tasks, aiming to assess LLMs’ abilities in understanding and meeting these constraints.
+Unlike LIFBench, these benchmarks only assess problem-solving capabilities in long-context scenarios, overlooking challenges of complex instruction following that arise in real-world applications.
 Given the complexities of evaluating instruction-following abilities, several studies introduce meaningful innovations in their evaluation methodologies to better tackle this multifaceted challenge.
@@
 (2024); Qin et al.
-(2024b); Zhang et al.
+(2024); Zhang et al.
 (2024a) decompose instructions into checklists composed of YES/NO questions or PASS/FAIL criteria, which answers should meet.
@@
 (2024) defines four criteria that should be assessed as they can encompass common errors made by LLMs, and develops automated evaluation metrics to reflect models’ ability in complex instruction-following scenarios.
-However, these studies fail to address model performance in long-context instruction-following scenarios and largely overlook stability in the instruction-following process.
-  Scenario Task ID #Samples
-List Single-ID LSI 180
-Multi-ID LMI 150
-Offset-ID LOI 396
-Offset-Element LOE 432
-Blur-ID LBI 396
-Blur-Element LBE 432
-MultiDoc Batch-label MB 150
-Find-dup-doc MF 150
-OneDoc Repeat OR 150
-QA OQ 180
-Extract OE 150
-Table 2: Statistics of LIFBench.
+However, these studies fail to concern evaluation in long-context instruction-following scenarios and largely overlook stability in the instruction-following process.
+Additionally, while the study Sakai et al.
+(2024) explores the impact of various prompt templates and languages on LLM, it focuses solely on NLU tasks.
+In summary, existing benchmarks either emphasize long-context scenarios or instruction-following capabilities, whereas LIFBench uniquely targets both simultaneously.
 3 LIFBench
@@
 The context $X$, as the main body of the prompt, provides essential information and varies by scenario.
-For example, in the List scenario, $X$ is a long list, while in the OneDoc scenario, $X$ represents a lengthy processed document.
+For example, in the List scenario (three scenarios are constructed in LIFBench, see Section 3.2), $X$ is a long list, while in the OneDoc scenario, $X$ represents a lengthy processed document.
 The parameter $len$ represents the number of tokens in $X$.
-The instruction $I$ is placed at the end of the prompt, which consists of two components: (1) the instruction template ($tpl$), which outlines the task requirements, and (2) the instruction variable ($var$), representing variable part within the template.
+The instruction $I$ is placed at the end of the prompt, which consists of two components: (1) the instruction template ($tpl$), outlining the task requirements, and (2) the instruction variable ($var$), representing variable part within the template.
 Generally, in LIFBench, $D$ and $I$ tend to be short, while $X$ is a long text with thousands of tokens.
 3.2 Dataset Construction
-To accurately simulate real-world application scenarios for LLMs in long-text processing, we construct 3 scenarios and 11 corresponding tasks (see Table 2), enabling a comprehensive and precise evaluation of instruction-following capabilities.
-List   The List scenario is designed to assess the ability of LLMs to handle structured batches of short-item information, as seen in applications like retrieval, recommendation, and structured data processing.
-The context $X$ is a lengthy ordered list, and LLMs must identify elements that meet the specified criteria provided in the instructions.
+To simulate real-world LLM applications in long-text processing, we construct 3 scenarios and 11 tasks (see Table 2) based on the following principles: (1) Task Diversity: Tasks should encompass varied constraints (e.g., format, quantity) to evaluate different instruction-following abilities; (2) Performance Distinguishability: Tasks must balance simplicity and complexity to distinguish model performance; (3) Input Scalability: Tasks should support extended input lengths to assess long-context capabilities; (4) Automated Evaluation: Task constraints can be assessed through an automated program.
+List   The List scenario tests how well LLMs can handle structured lists, such as retrieving specific items and processing structured data.
+The input $X$ is an ordered list with UUIDs and natural language instructions.
 The scenario includes six tasks: Single-ID and Multi-ID focus on retrieving specific elements from an ordered list using provided IDs.
@@
 "ID" and "Element" refer to the type of reference in the retrieval process, representing either the position number in the ordered list or the list element itself.
-MultiDoc   The MultiDoc scenario tests LLMs on tasks involving coarser-grained information, such as multi-document summarization, clustering, and retrieval-augmented generation (RAG).
-Unlike the List scenario, the MultiDoc tasks require models to compare multiple documents, identify differences, and execute batch operations, demanding holistic attention to the input rather than focusing on a few key elements.
-To construct context $X$, we curate and process datasets from diverse domains, combining them into multi-document collections with high diversity.
-There are two tasks in this scenario.
-Find-dup-doc requires LLMs to identify and output multiple duplicate documents in a specified format.
-In batch-label, models label each document based on specified attributes.
-OneDoc   The OneDoc scenario simulates tasks involving processing of a single and very long document, typically occurring in use cases like information extraction and question answering.
-In this scenario, LLMs are required to identify and process the most relevant information from a vast amount of text, testing their ability to resist noise and focus on task-specific instructions.
-To create this scenario, we construct an extra-long document by concatenating entries from the Paul Graham Essays dataset [^2] https://huggingface.co/datasets/sgoel9/paul_graham_essays, and tagging some sentences as key information.
-Then three tasks are designed: The Repeat task requires models to repeat a specified amount of key information.
-The Extract task instructs models to extract specific types of key information.
-The QA task presents models with a sentence and asks it to identify whether the sentence is labeled as key information.
+MultiDoc   The MultiDoc scenario evaluates how well LLMs process multiple documents, such as summarization and retrieval.
+Models need to compare documents, find differences, and handle batch operations.
+The input $X$ consists of multi-document collections from diverse sources.
+Each document has six fields: "text", "id", "iD2", "title", "date", and "source", with a length of 300–500 tokens.
+Tasks include Find-dup-doc, which finds duplicate documents, and Batch-label, which assigns labels to documents based on given attributes.
+OneDoc   The OneDoc scenario tests how well LLMs process a single long document, such as extracting key information or answering questions.
+A long document is created by combining essays from the Paul Graham Essays dataset [^2] https://huggingface.co/datasets/sgoel9/paul_graham_essays.
+Some sentences are marked as key information.
+Tasks include Repeat, where models repeat a given amount of key information; Extract, where models extract specific key details; and QA, where models check if a sentence is labeled as key information.
 We manually write scenario descriptions and instruction templates for all tasks, with detailed examples provided in Appendix H.
@@
 Specifically, we modify the number of elements in the List or the number of documents in MultiDoc and OneDoc to achieve the desired length.
-We have pre-constructed ample corpus for each scenario, supporting expansions up to 2M tokens in one prompt.
-Expression   In real-world contexts, due to personality and individual differences, individuals often provide significantly varied descriptions of the same subject, which undoubtedly introduces challenges to the stability of large models in following instructions.
+Ample corpus are pre-constructed for each scenario, supporting expansions up to 2M tokens in one prompt.
+Expression   In real-world contexts, due to personality and individual differences, individuals often provide significantly varied descriptions of the same subject, which undoubtedly challenges the stability of large models in following instructions.
 To assess LLM robustness in this regard, we diversify instruction templates to create multiple expressions style with differing wording and syntax.
 Our approach follows a four-step process, namely "Rewriting-Encoding-Clustering-Sampling (RECS)".
-First, we use GPT-4 Achiam et al.
-(2023) and Claude Anthropic (2024) to generate 40+ rewrites of each original instruction template.
+First, to ensure diversity and mitigate biases from any single model, we use GPT-4 Achiam et al.
+(2023) and Claude Anthropic (2024) to generate 40+ rewrites of each original instruction template, complemented by a subsequent manual review to further validate and refine the outputs.
 Next, the rewritten templates are encoded into vector representations for clustering, with the number of clusters set to the target number of rewritten instructions for each task.
 For instance, if a task needs five rewritten instructions, we will create five clusters.
-Finally, we sample from each cluster, selecting the usable template closest to the cluster center as the final diversified expression.
-More implementation details and evidence of effectiveness can be found in Appendix B.
+Finally, from each cluster, we select the usable template nearest to the center as the final diversified expression.
+Further details and effectiveness evidence are provided in Appendix B.
+  Scenario Task ID #Exp. #Var. #Data
+List Single-ID LSI 5 6 180
+Multi-ID LMI 5 5 150
+Offset-ID LOI 11 6 396
+Offset-Element LOE 12 6 432
+Blur-ID LBI 11 6 396
+Blur-Element LBE 12 6 432
+MultiDoc Batch-label MB 5 5 150
+Find-dup-doc MF 5 5 150
+OneDoc Repeat OR 5 5 150
+QA OQ 5 6 180
+Extract OE 5 5 150
+Table 2: Statistics of LIFBench. #Exp. and #Var. represent the count in the data extension of Expression and Variable.
 Variable   Some placeholders are preset in the instruction templates to indicate variable parts of instructions, i.e., instruction variables.
 These variables encompass elements such as query keys (for retrieval tasks), categorization criteria (for classification tasks), and format requirements.
+For position-related or numerical variables, we maintain an even distribution, and manually adjust other variables during task iterations to precisely control the task difficulty (see examples in Appendix C).
 By analyzing LLM performance across varying instruction variables, we can evaluate the models’ understanding and consistency in executing instructions.
 Ideally, models with strong instruction-following abilities execute instructions stably across variable conditions, while weaker models may exhibit inconsistency.
-Each instruction variable has been manually assigned by specific sampling spaces and rules, enabling the generation of a comprehensive instruction set.
-For specific examples, please refer to Appendix C.
 Table 2 details the task counts for each scenario. Compared to the other two scenarios, tasks in the List scenario are simpler, so they carry less weight (see Section 4.1) but are more numerous.
-Figure 2: Rubric score distribution across different scenarios and capabilities.
-We carefully designed the rubric to ensure scores are as evenly distributed as possible across different scenarios and capabilities.
+Figure 2: Rubric score distribution across different scenarios and capabilities. Scores across different scenarios and capabilities tend to be evenly distributed.
 4 LIFEval
@@
 To evaluate the output quality of LLMs on LIFBench tasks, we introduce Automated Rubric-based Scoring (ARS), an accurate and efficient programmatic evaluation method.
-As shown in Figure 1 and Figure 2, we manually craft scoring rubrics $\mathcal{R}$ for each task.
-Each rubric consists of several scoring points $s$ and is assigned a weight $\tilde{s}$ according to its importance and complexity.
-In other words, a larger $\tilde{s}$ means a more complex task, requiring more evaluative consideration.
+As shown in Table 6 in Appendix D, we manually craft scoring rubrics $\mathcal{R}$ for each task.
+Each rubric consists of several scoring points $s$ and is assigned a weight $\tilde{s}$ according to its complexity and difficulty.
+In other words, a larger $\tilde{s}$ means greater complexity, requiring more steps for evaluation.
 All of these points can be assessed automatically through a program.
@@
 In this equation, $\mathcal{A}_{t}$ represents the model’s responses on task $t$, and $\tilde{R}_{t}$ is the sum of weights across all scoring points.
-The function $f_{s}(\cdot)\in[0,\tilde{s}]$ represents the average score of all outputs for scoring point $s$, as determined by the automated evaluation program.
+The function $f_{s}(\cdot)\in[0,\tilde{s}]$ represents the average score of all outputs for scoring point $s$.
+Its implementation relies on programmatic evaluation pipelines tailored to each scoring point.
+For example, in structural verification (e.g., JSON Dict format, see Program 1), we use a multi-stage detection process that includes symbol check, Parsing check, and KV Check.
+These stages progressively impose stricter requirements on the model’s output, with final scores based on performance at each stage.
+Overall, while the core logic is manually engineered according to the rubric, the execution is fully automated through programs.
+We provide detailed considerations regarding the scoring rubric, as well as additional examples of evaluation programs, in Appendix D.
 Naturally, defining $\mathcal{T}=\{t_{i}\}_{i=1}^{N_{t}}$ as the set of all tasks, the overall test result for any model on LIFBench is the weighted average of scores across $\mathcal{T}$:
 $ARS_{overall}=\frac{\sum_{t\in\mathcal{T}}\tilde{R}_{t}\cdot ARS_{t}}{\sum_{t\in\mathcal{T}}\tilde{R}_{t}}\,.$ (3)
+Although a well-designed scoring rubric enables LIFEval to provide a more comprehensive and rigorous evaluation, the relatively high human cost inevitably limits its generalizability.
+To address this, when extending LIFBench with new samples or adapting LIFEval to other datasets, users may adopt a simplified approach by replacing weighted averaging in Equation 2 and 3 with a simple mean.
+This streamlined method bypasses the rubric design process while still satisfying the requirements of many practical applications.
   OneDoc List MultiDoc
@@
 where $\mathbb{I}(s,c)$ is the indicator function that identifies whether a scoring point $s$ is related to the capability $c$.
-We manually crafte this mapping so that when a scoring point $s$ is relevant to the capability $c$, the indicator $\mathbb{I}(s,c)$ equals 1, and 0 otherwise.
+We manually crafted this mapping so that when a scoring point $s$ is relevant to the capability $c$, the indicator $\mathbb{I}(s,c)$ equals 1, and 0 otherwise.
+Table 6 provides more detailed information on this.
+Importantly, as shown in Figure 2, scores across different scenarios and capabilities are carefully balanced during data collection to ensure less bias in LIFEval.
 4.3 Instruction Following Stability
@@
 $\begin{split}\mathcal{Y}_{t}^{p}=\{ARS_{t}(\mathcal{A}_{t}^{j})\}_{j=0}^{N_{p}}\,.\end{split}$ (5)
-Finally, the Instruction Following Stability (IFS) on task $t$ is calculated as the variance (denoted by $\sigma(\cdot)$) of the performance values $Y$ divided by their mean, formally expressed as:
+Finally, the Instruction Following Stability (IFS) on task $t$ is calculated as the standard deviation (denoted by $\sigma(\cdot)$) of the performance values $Y$ divided by their mean, formally expressed as:
 $\begin{split}IFS_{t}^{p}=\frac{\sigma(\mathcal{Y}_{t}^{p})}{\bar{\mathcal{Y}}_{t}^{p}}\in[0,+\infty)\end{split}$ (6)
@@
 (2023).
-Experiments are conducted across six context length intervals from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation.
+The experiments were conducted across six context length intervals, ranging from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation.
 Token counts are calculated using GPT-4’s tokenizer [^3] https://github.com/openai/tiktoken, and truncation is applied to adjust context $X$ for models unable to process the longest contexts.
-Further details about the experimental setup can be found in the Appendix F.
+In data extension, 5–6 template variants are created for each original instruction, and each instruction variable has 5–10 candidates for sampling.
+Further details can be found in Appendix E.
   IFS Overall
@@
 […]
-Table 4: Instruction Following Stability from different perspectives.
-The integers in parentheses represent rankings, where smaller numbers indicate better performance.
-The overall order in each partition depends on the average IFS rankings from three perspectives.
-$\dagger$ indicates that the context $X$ on the longest interval is right-truncated.
+Table 4: Instruction Following Stability (IFS) from three perspectives, with rankings in parentheses (smaller is better).
+The overall order is based on average IFS rankings.
+$\dagger$ denotes truncation of context $X$ on the longest interval.
 5.2 Results on LIFBench
@@
 While closed-source models dominate most tasks, open-source models excel in only a few (e.g., Blur-ID and Batch-label).
-Overall, open-source small and medium models still lag behind closed-source large models.
+Moreover, compared to LSI, most of models perform worse on LOI and LBI tasks, likely due to LSI’s closer alignment with their training data, highlighting the need to enhance their ability to handle complex instructions.
+Similarly, the model performance on LBE (LOE) tasks is better than that on LBI (LOI), indicating that referencing specific elements is easier than using non-semantic IDs.
 Capability-categorized Performance   As shown in Figure 3, closed-source models lead in all dimensions.
 Within each model series, the solid lines consistently enclose their corresponding dashed lines, highlighting the significant impact of instruction fine-tuning on overall performance.
-Notably, the Format dimension exhibits the most compact clustering, suggesting that formatting requirements are well-covered across training corpora for most models.
+Notably, the Format dimension shows the tightest clustering, suggesting that formatting requirements are well-covered across training corpora for most models.
 In contrast, the Recog dimension shows marked differences, likely reflecting varying levels of effort in incorporating numerical cognition data during training.
+Figure 3: The Instruction Following Performance in six capabilities.
+Lines of the same color in the chart represent models from the same series.
+Dashed lines represent base models, while solid lines represent their officially fine-tuned variants.
+Figure 4: Overall ARS score (a) and instruction following stability (b) under different input length.
 Stability   Table 4 presents the IFS scores and rankings from three perspectives, revealing discrepancies between model stability and task completion ability.
@@
 surpassed all closed-source models in "Expression".
-Figure 3: The Instruction Following Performance in six capabilities.
-Lines of the same color in the chart represent models from the same series.
-Dashed lines represent base models, while solid lines represent their officially fine-tuned variants.
-Figure 4: Overall ARS score (a) and instruction following stability (b) under different input length.
 Figure 5: Instruction following performance in six core capabilities under different input length.
+Figure 6: The impact of increasing context length on ARS scores in tasks with different instruction complexities.
 5.3 Effects of longer inputs
@@
 experiences a sharp drop, indicating its limited ability to handle extended inputs.
-The negative impact of input length is also evident in stability metrics.
+The negative impact of input length is evident in stability metrics.
 As illustrated in Figure 4(b), most models demonstrate poorer stability in long-context scenarios, with the lowest stability observed at the longest input lengths.
@@
 Conversely, "Recog" experiences the steepest decline, highlighting the challenges models face in maintaining recognition ability as input length grows.
-Model size also plays a crucial role in long-context instruction-following performance.
+Model size also plays a crucial role in long-context instruction-following.
 Larger models generally are better at capabilities like "Ori," "Num," and "Spat," indicating that handling original content and spatial reasoning demands more model parameters.
 In contrast, format-related abilities are effectively managed even by smaller models, suggesting they are less dependent on model size.
+5.4 Impact of Instruction Complexity
+To analyze how context length and instruction complexity simultaneously affect model performance, we divide tasks into two groups based on scoring weight: easy ($\leq$10) and hard (>10).
+As noted in Section 4.1, higher weights indicate greater challenges, with 10 chosen to balance the number of tasks in each group.
+As shown in Figure 6, the performance of different models varies significantly.
+Llama-3.1-70B-Inst.
+struggles with long contexts across all tasks, highlighting its difficulty in handling long inputs.
+For GPT-4 and Qwen2.5-72B-Inst., the performance on both groups is similar in short contexts, but has sharper declines on hard tasks with longer contexts, revealing their limitations in handling complex tasks in long-context scenarios.
+Other models exhibit larger performance gaps even in short input (i.e., 4k tokens), indicating greater task difficulty impact for them.
+Generally, hard group degrades more significantly as context length increases, suggesting a compounded negative effect of task complexity and context length on LLM performance.
 6 Conclusion
@@
 We leave the automation or reduction of these manual efforts for future work.
+Acknowledgments
+This work was supported in part by National Natural Science Foundation of China (No.
+92270119) and Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education.
 Appendix A Details in Data Collection
@@
 Implementation details   To implement RECS for expression extension, we chose the BGE-M3 text embedding model Chen et al.
-(2024) for encoding and applied K-means clustering Macqueen (1967).
-Additionally, since LLMs can introduce inaccuracies in rewrites due to misinterpretations, we manually filtered out unsuitable rewrites during the sampling phase, resulting in 5~10 instruction templates for each task.
+(2024a) for encoding and applied K-means clustering Macqueen (1967).
+Additionally, since LLMs can introduce inaccuracies in rewrites due to misinterpretations, we manually filtered out unsuitable rewrites during the sampling phase, resulting in 5~6 instruction templates for each task.
 Prompt for Rewriting in RECS Please rewrite the given prompt according to the following requirements: 1.
@@
 Specifically, after the Rewriting phase, we constructed a dataset by randomly sampling (rather than clustering) the same number of instructions and calculated the IFS values for sixteen models on this dataset.
-As shown in 6, compared to random sampling, RECS led to higher expression IFS scores in 75% of the models, outperforming the IFS improvement rates observed for the Variable and Length perspectives, at 37.5% and 62.5% respectively.
+As shown in 7, compared to random sampling, RECS led to higher expression IFS scores in 75% of the models, outperforming the IFS improvement rates observed for the Variable and Length perspectives, at 37.5% and 62.5% respectively.
 Additionally, when comparing the mean IFS values across all models, we observed that RECS improved expression IFS while slightly lowering variable IFS and length IFS.
@@
 Appendix C Sampling Space for Instruction Variables
-As shown in Figure 8, we reserved identical placeholders for each instruction template and set the sampling space accordingly.
+As shown in Figure 9, we reserve identical placeholders for each instruction template and set the sampling space accordingly.
 Depending on the task requirements, the types of instruction variables include numerical values, lists, phrases, long sentences, and format indicators, etc.
+We carefully consider variable distribution to avoid biases from the sampling mechanism.
+For example, in position-related variables within the List scenario, inputs were divided into three sections (beginning, middle, and end), with the middle section receiving the largest portion.
+We then randomly sample 2–3 elements from each section to ensure balance.
+Additionally, we introduce unconventional causal relationships in certain rule-based variables to increase task difficulty.
+For instance, in the QA task, models are required to use "False" for correct answers and "True" for incorrect ones, making it more challenging to follow instructions.
 List: Single-ID Instruction templates: 1. Retrieve the entry at position {pos} in the list. Display it immediately. 2. What is the {pos} element in the list? Output it directly. 3. … Sampling Space
@@
 Figure 9: Examples of the sampling space. Red indicates the instruction variables $var$, and {} indicates a placeholder in the instruction.
-Appendix D Rubric Design and Program Evaluation
-Rubric Design   As shown in Table 5 and Figure 2, we manually designed a scoring rubric for LIFBench, ensuring a balanced evaluation across different scenarios and capabilities.
-The rubric assigns greater weight $\tilde{s}$ to score point $s$ that is more challenging.
-For instance, in assessing format correctness, compared to the QA (OQ) task, the Batch-Label (MB) task are assigned higher scoring weights due to more complex formatting requirements.
-Program Evaluation   As described in Section 4.1, we design an automated evaluation program based on the scoring rubric to evaluate the quality of LLM responses.
+Appendix D Rubric Design and Evaluation Program
+Rubric Design   For the weight $\tilde{s}$ assignment of scoring points, we primarily considered three factors:
+(1) Evaluation complexity: Scoring points with higher evaluation complexity require more assessment steps and are therefore assigned greater weight.
+For example, in assessing format correctness, the Batch-Label (MB) task is given higher scoring weights compared to the QA (OQ) task due to its more complex formatting requirements.
+(2) Task difficulty for models: Some scoring points may not be as complex to evaluate but are particularly challenging for the model to fulfill.
+In such cases, we allocate greater weight to these points as well.
+(3) Balance across capabilities and scenarios: Based on the considerations in (1) and (2), we made fine adjustments to ensure that the weights assigned to different instruction-following capabilities and scenarios are as balanced as possible.
+This balance is shown in Figure 2.
+On this basis, we present the rubric of LIFEval in Table 6.
+Additionally, the pseudocode examples reflect our considerations above.
+Program 1 show the correspondence between evaluation steps and their respective weights.
+In Program 2, we award an additional point for responses that fully meet the requirements, highlighting the difficulty of that particular scoring point.
+Evaluation Program   As described in Section 4.1, we design an automated evaluation program based on the scoring rubric to evaluate the quality of LLM responses.
 Inspired by Zhou et al.
@@
 However, LLM responses are not always well-structured and may contain minor errors that hinder the extraction of structured information.
-To mitigate this, we carefully designed task-specific validation programs that employ techniques such as regex matching and substring retrieval to maximize the extraction of valid part from LLM responses.
-Finally, to ensure evaluation quality, we manually provided reference answers for all examples in LIFBench, verifying that they can achieve full scores in our automated evaluation process.
-Appendix E Inspiration for IFS
+To mitigate this, we carefully designed task-specific validation programs that employ techniques such as regex matching and substring retrieval to maximize the extraction of valid part from LLM responses (see Program 1).
+Finally, we manually provided reference answers for all samples in LIFBench.
+We verified that the automated evaluation programs in LIFEval consistently achieve full scores on all reference answers, thereby fulfilling the first principle outlined above.
+Appendix E Experiments Setup
+Baselines   For closed-source models, we specifically use gpt-4o-2024-08-06 and gpt-4-0125-preview to represent the performance of GPT-4o and GPT-4, respectively.
+For open-source models, all implementations were sourced from Hugging Face [^5] https://huggingface.co/ (as detailed in Table 5), and we test the performance of both base models and fine-tuned models.
+Models with the suffix "-Inst." or "-chat" indicate that they have been fine-tuned on instruction or dialogue data.
+Inference   During the inference process, we complete the deployment of all open-source models with the vLLM Kwon et al.
+(2023).
+The temperature is set to 0 to ensure deterministic outputs.
+For the SFT versions of the models, we used the official chat template.
+For the base versions, we add a suffix "Output: " to prompt the models to generate answers according to the instructions.
+Token counts were calculated using GPT-4’s tokenizer [^6] https://github.com/openai/tiktoken, and truncation was applied to adjust context for models unable to process the longest contexts.
+To ensure diversity and accuracy in instruction phrasing, we generated 5–6 template variants for each original instruction and manually filtered them.
+Each template included placeholders for variables (e.g., numerical values, lists, phrases, long sentences, and format indicators), dynamically sampled from curated sets of 5–10 candidates to create task-specific prompts.
+In order to avoid unnecessary time consumption resulting from endless repetitions in the output, we set maximum generation lengths for different tasks: List scenario tasks that output a single element were limited to 100 tokens, tasks in the MultiDoc scenario are set to 4096 tokens, and the rest of the tasks are limited to 512 tokens.
+Due to variations in encoding efficiency across models’ tokenizer, some models can not accommodate the longest inputs.
+To address this, we apply right truncation to the context $X$, ensuring the completeness of scenario description $D$ and instruction $I$.
+The experiments were conducted across six context length intervals, ranging from 4k to 128k tokens, with task-specific output limits to ensure sufficient space for model generation.
+  Model Hugging Face ID
+Llama-3.1-70B-Inst. meta-llama/Llama-3.1-70B-Instruct
+Llama-3.1-8B-Inst. meta-llama/Llama-3.1-8B-Instruct
+Llama-3.1-70B meta-llama/Llama-3.1-70B
+Llama-3.1-8B meta-llama/Llama-3.1-8B
+[…]
+Qwen2.5-32B-Inst. Qwen/Qwen2.5-72B-Instruct
+Qwen2.5-14B-Inst. Qwen/Qwen2.5-72B-Instruct
+Qwen2.5-7B-Inst. Qwen/Qwen2.5-72B-Instruct
+Qwen2.5-72B Qwen/Qwen2.5-72B
+Qwen2.5-32B Qwen/Qwen2.5-72B
+Qwen2.5-14B Qwen/Qwen2.5-72B
+Qwen2.5-7B Qwen/Qwen2.5-72B
+[…]
+C4AI-cmd-r-v01 CohereLabs/c4ai-command-r-v01
+[…]
+LWM-Text-1M LargeWorldModel/LWM-Text-1M
+[…]
+[…]
+Table 5: A mapping between the Hugging Face model IDs and the aliases used in the paper.
+Appendix F Inspiration for IFS
 IFS is derived from the coefficient of variation (CV) Brown (1998), a statistical measure widely used in fields such as risk assessment and quality control Shechtman (2013).
@@
 By capturing fluctuations in model performance relative to its average capability, IFS offers a robust and fair metric for comparing the stability of different LLMs.
-Appendix F Experiments Setup
-Baselines   For closed-source models, we specifically use gpt-4o-2024-08-06 and gpt-4-0125-preview to represent the performance of GPT-4o and GPT-4, respectively.
-For open-source models, we test the performance of both base models and fine-tuned models.
-Models with the suffix "-Inst." or "-chat" indicate that they have been fine-tuned on instruction or dialogue data.
-Inference   During the inference process, we complete the deployment of all open-source models with the vLLM Kwon et al.
-(2023).
-The temperature is set to 0 to ensure deterministic outputs.
-For the SFT versions of the models, we used the official chat template.
-For the base versions, we add a suffix "Output: " to prompt the models to generate answers according to the instructions.
-In order to avoid unnecessary time consumption resulting from endless repetitions in the output, we set maximum generation lengths for different tasks: List scenario tasks that output a single element were limited to 100 tokens, tasks in the MultiDoc scenario are set to 4096 tokens, and the rest of the tasks are limited to 512 tokens.
-Due to variations in encoding efficiency across models’ tokenizer, some models can not accommodate the longest inputs.
-To address this, we apply right truncation to the context $X$, ensuring the completeness of scenario description $D$ and instruction $I$.
+Program 1 Format Validation (JSON Dict)
+Target: Verify if the output meets the JSON Dict format.
+1:LLM answer $a$, full score $\tilde{s}=4$
+2:Validation score $s_{a}=f_{s}(a)\in[0,\tilde{s}]$
+3:Step 1: Symbol Check (Max 1 point)
+4:$s_{a}\leftarrow 0$
+[…]
+6:     $s_{a}\leftarrow s_{a}+1$
+7:end if
+8:Step 2: Parsing Check (Max 2 points)
+[…]
+[…]
+11:     $s_{a}\leftarrow s_{a}+2$
+[…]
+13:     $s_{a}\leftarrow s_{a}+1$
+14:end if
+15:Step 3: KV Check (Max 1 point)
+[…]
+17:     $s_{a}\leftarrow s_{a}+1$
+18:end if
+19:return $s_{a}$
+[…]
+Program 2 Quantity Verification (LMI)
+Target: Verify if the number of output elements is correct.
+1:LLM answer $a$, target number $n_{\text{gold}}$, full score $\tilde{s}=3$
+2:Validation score $s_{a}=f_{s}(a)\in[0,\tilde{s}]$
+3:$n_{\text{pred}}\leftarrow\text{extract\_quantity}(a)$
+4$\triangleright$ Extract the number of elements in the output
+5:if $n_{\text{pred}}=n_{\text{gold}}$ then
+6:     $s_{a}\leftarrow\tilde{s}$triangleright Full score if the quantity matches
+7:else
+8:     $s_{a}\leftarrow\max\left(0,\left(1.0-\frac{|n_{\text{pred}}-n_{\text{gold}}|}{n_{\text{gold}}}\right)\right)\times 2$
+9$\triangleright$ Penalty based on relative deviation
+10:end if
+11:return $s_{a}$
+Note: The function $n_{\text{pred}}=\text{extract\_quantity}(a)$ can be executed in various ways.
+If the LLM answer $a$ passes format checks and is parsable, the target quantity is directly obtained.
+Otherwise, string processing techniques are employed to extract it with feedback.
   Tasks Score Point $s$ Weight $\tilde{s}$ Related Capability
@@
 Order correctness. 2 Spat
-Number of output elements is correct. 3 Num
+The number of output elements is correct. 3 Num
 Answers correctness. 3 Ori